A Bibliography of Publications in British Society for the History of Mathematics Newsletter and BSHM Bulletin: Journal of the British Society for the

History of Mathematics and British Journal for the History of Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

03 August 2024 Version 1.18

Title word cross-reference

$15.00 [Cla20]. $24.95 [Ran21]. ϵ − δ [Nak14]. L [Coa08].

-functions [Coa08].

0 [Ran21].

1 [Rog20, Wil17a]. 1240s [Big21]. 138pp [Cla20]. 1500s [Nas16, Nas16]. 1680s [Ano23h, Ano24a]. 17 [Cra16]. 17th [Rob18, Wes16]. 1850 [SS22]. 1925 [Fin19].

2 [Wil17b]. 2- [Hol06]. 2006 [Bur06]. 2015 [Cri16, Flo16]. 2016 [Hol16b, Led16a, Led16b, McC16, PR18, Wes16]. 2019 [Ano19a]. 2023 [Ano23m, Ano24d]. 29th [Ste19].

3 [Sch21].

4 [Fin19]. 4- [Hol06]. 400th [Sim14]. 448pp [Dau20]. 492pp [Hit21].

5 [Bel15b, Mac21].

6 [Hit21].

7 [Cri20a, McQ20].

8-square [Hol06].

9 [Cla20, Dau20]. 978 [Cla20, Cri20a, Dau20, Fin19, Hit21, Mac21, McQ20, Ran21, Sch21]. 978-0-19-885224-7 [Cri20a]. 978-0-19027-185-5 [Mac21]. 978-0-1987-3311-9 [Dau20]. 978-0-6911-9172-0 [Ran21]. 978-1-4704-4889-9 [Cla20]. 978-1-64327-189-7 [McQ20]. 978-1-78205-290-6 [Hit21]. 978-2-7288-0603-4 [Fin19]. 978-3-03001-071-3 [Sch21]. 99 [Tan20]. 9th [RK19].

Abeles [Law18]. abelian [Sør13, Sør13]. abstract [Hol14]. Academician [Hol18b]. account [BBE^{+}21, Man10]. accounting [CR08]. accuracy [Row24]. Acheson [Fea18]. Achilles [Law18]. acquisition [Sat21]. activities [Zel19]. activity [Rek19, Row15]. actuary [Big22]. Ada [Flo16, HMR17]. added [Boy24]. Addendum [Ano24a, GG14a]. addresses} [Cra12]. adornment [dB09]. adult [Kay06]. advancement [Bee17]. adventure [Bel10, Fea18, Law16]. Africa [Mac20]. African [Cla20]. after [Rod18]. age [Wes22, Daw18]. AI [Row24]. air [BRO24]. al [Hol18a]. Alex [Fal20b]. Alexander [Car18, Fal20b]. algebra [Ano23o, BK22, Fre13, Gag17, Hol16a, Jan14, Rau16, Ric16]. algebraic [Ano23o, Par17]. algebras [Par17]. Alice [Ken12, vW22]. alterations [Bee20]. Altro [Gui18]. altum [Cra17]. Amatino [Smi08]. America [Den15, Cam23]. American [Cla20, Abr20, DeC16, Kas07, Abr18]. Americans [Zel19]. Amirouche [Law18]. Analyse [Bra13]. analyses [Ala20]. Analysis [YO17, ADD16, BK22, Bur06, EMSS22]. analyticae [Sel08]. analytical [Ano23i, Cra20b, Pad17]. anamorphic [HS15]. Anastácio [Dom14]. Ancient [Car18, CR08, Cla10, Cuo12]. Andrew [Ric14, War07]. Andrews [Ams08]. anecdote [Ano23n]. Anglo [PR18]. Anniversary [Sim14]. Annual [CK07]. Anstice [AC13b]. antecedents [BBG^{+}20]. Anti [Hol15]. Anti-Semitism [Hol15]. anticolonial [Ano23f]. Antikythera [Car18]. antiquity [Nas16, Ric16]. Anton [Hol09]. Apollonius [Kun07]. apologia [Cra17]. appearance [Bra15]. application [Ano23a, Rei23]. applied [MG11, Rau16]. appreciation [GG14b]. approach [Pet15, Rog20]. approaches [Sør17]. appropriations [Sør17]. approximation [Man23]. April [Poo15]. aqua [Cra17]. Arabic [Oak22]. Arbuthnot [DB21]. arc [Ano23e]. Archimedes [Bur06, Lee13]. Architecture [Nas16, Nas16]. archive [Ald21]. area [Atz15, Mag12]. argument [Mok18]. Arianrhod [Mac21, Opi16]. arithmetic [Big09, Cag19, Mas07, Oak22, Wes12]. Arithmetical [Den06]. arithmetization [Hol16a]. art [Gra20]. Arte [RVME18]. Arthur [Cra06]. article [Ano07d]. Artin [Sch15]. Artis [Sel08]. arts [PO13]. Ashley [Sat21]. associated [CO11]. Asteroid [Bus18]. Astronomer [Rob18]. Astronomia [Rec21]. Astronomy [Dau20]. asymmetry [Neu08]. Atlas [Hil14]. Aubin [Fin19]. Audin [HG12]. August [Emi12, PR18]. Augusto [MS22]. authority [Dau20]. automation [MG11]. Avoiding [Wes12, Nak14]. Award [Man10]. Awards [Ano24b, Ano23k]. away [ACO19].

B [NB21, NB21]. Babbage [Bow15, Pad17]. ‘Baby’ [And09]. Bache [Ken12]. back [Mac20, Rau13]. Bad [Man10]. Bagatelle [Pri06]. Bails [Ano23i]. Ball [McD18]. ballistic [Bar09]. balls [Man12]. Barba [NF23]. Baron [Cra12]. barriers [Tha10]. Bartering [Mas07]. based [Ano23e, Law06, Rog20]. Bath [PR18]. Batterson [Abr18]. BCE [SS22]. be [Emi12]. Beating [Cri20b]. Beaud [Sta15]. beautiful [Woo19]. beauty [KS17]. became [Ast20, Big19, Cri18b]. Beeley [Poo15]. began [Bow15]. beginnings [Fri16]. Belfast [Tur15]. bell [Man10, Hol16a]. Benito [Ano23i]. Benjamin [Bar09, Wes24]. Berlinski [Sia15]. Bernardi [Rob18]. Bernoulli [Ala20, Bra13]. Bernoullian [Tou19]. Bernt [Chr09]. Berry [Sen16]. Bertrand [Gau10]. Between [Ano23a, PO13, BBG^{+}20, Bru20, Fri19, Pis16, Pri08, Prz09, Rei23, Rod17b, Tay11]. beyond [HG12, Hol20, McD18, Sen16]. Bibby [Bur06]. Bicentenary [Flo16, AC13b]. bigger [SW17]. Biography [Hol17, Pad17, Sch21]. biology [Led19]. Birkbeck [Hol16b]. Birkhäuser [Sch21]. Birmingham [Cri16]. birth [Spe19, Law16]. bisection [Rec21]. blanks [vB10]. blind [BWH22]. Blundeville [Son17]. Board [Ano14c, Ano15]. Bolzano [BBG^{+}20]. Book [Abr18, Ano23a, Big19, Bow15, Bus18, Cal18, Car18, Cla20, Cra15, Cri18a, Cri20a, Dau20, Daw18, Den15, Fea18, Fin19, Hig19, Hit21, Hol15, Hol18a, Hol18b, Jes15, KJB11, Law16, Mac21, Man15a, McD18, McQ20, Nas16, Opi16, Pec20, Poo15, Ran21, Ric14, Ric16, Rob18, Rod18, Rog20, Sch21, Sch15, Sen16, Sia15, Sim18, Son14, Spe19, Sta15, Tan20, Van19, Wag18, Wil18a, Woo19, Bas20, Sat21]. booklet [Rod17a]. Books [Ano23l, Seg20, Agg07, Law11, Mas07, Sat21, Den15]. Boole [Mok18, Mur16, Sim08, Hit21]. Boolean [Jan14]. Boran [Hig19]. Boris [Hol18b]. Bouk [Big19]. bounded [PCMA15]. Brahmagupta [Atz15]. braid [Fri19]. Brain [Row14, CK07]. Branch [Wil15e]. Brass [Row14]. Breaking [Tha10]. breakthroughs [Cra16]. Brechenmacher [Hol18a]. Brian [Cri18a]. bridges [Mal08]. brief [McD18, Per15]. Briggs [Jag06, Son17]. bringing [vB10]. Britain [DB21, Dom14]. British [AH08, Ald21, Ano19e, Ano23b, Ano24b, Dav18, Gra06, Par06, Ric06, Sim14, Wil15a]. Brown [Jes15]. Brummelen [KJB11]. BSHM [CK07, Flo13, Man15c, Ric15a, Wil15g, Ano06a, Ano06b, Ano06c, Ano07a, Ano07b, Ano07c, Ano08a, Ano08b, Ano08c, Ano09a, Ano09b, Ano10a, Ano10b, Ano11a, Ano11b, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, Ano14b, Ano18a, Ano19b, Ano19c, Ano20a, Ano20b, Ano21a, Ano21b, Ano22a, Ano22b, Ano22c, Ano23c, Ano23d, Ano23b, Ano24c, Ano24d, Cri15, Cri16, Fal18, FC20, Flo16, Ram11b, SPR09, SW12, Ste23, Tur15, Wal14, Wee10, Wes18, Wil13, Wil15b, Wil15e, Wil18b, WF21]. BSHM/Gresham [Ric15a, Flo16]. BSHM/Institute [Tur15]. BSHM/Lewis [Wil18b]. Bulletin [Ano05a, Ano05b, Ano05c]. bush [Rau13].

C [Jes15, Sch21, Ver16]. C. [Wil18b]. Calculation [Fro15]. Calculus [Fea18, Ote21, BBE^{+}21, Ber21, Cal18, Jes15, Nak14]. Calinger [McQ20, Wil18a]. Cambridge [Cra20b]. came [Rau13]. campaign [BRO24]. Canadian [Man15c]. Canonis [Sim14]. cap [Har09]. capacity [MM19]. Capecchi [Spe19]. Caramuel [CRPH19]. Career [Ano19a, Ano20d]. Carroll [Law18, Pec20, Wil18b, Law18, Mok18]. Carrollian [Law18]. cartoonist [Pad17]. case [Fad15, Fri19, Gra06, Hol18b, Jes15, RB16a, RB16b, Zel19]. Cassini [Rob18]. catalogues [Mok18, Sat21]. catching [Abr18]. Cauchy [BBG^{+}20]. Cédric [Law16]. Celebration [Flo16, Hol17, Sim14]. central [Den15]. Century [Wes16, Agg07, Ano23b, Big09, Big17, Big22, BK22, Dec11, Dom14, Edw06, Gag17, Gra06, Gul22, Jen10, Kar07, Kle07, KS15, Krü10, Lor12, RVME18, RK19, Seg20, YO17, Zel19, dB09, Ano18a, Rob18, Ric16]. Cerdà [Ber21]. Ceres [Bus18]. Châtelet [Opi16]. chalk [TCP19]. chance [Pri08, Cri18a]. chancery [Big21]. changed [Cra16, Sim18]. character [Nak14]. Charles [Pec20, War17]. charm [Ast20]. charts [MG11]. chicken [Rau13]. Children [Cra06]. China [Dau20]. chorograph [Cra10]. Chris [Ano07d]. Christmas [Cri16, Wil15b]. Christoph [Poo15, GG14b]. Christopher [Dau20]. Chronicle [Sen16]. ciphers [Not20]. circle [Bee20]. circuits [Jan14]. cities [Min09]. city [dB09]. claim [Hol14]. classmate [Cri18b]. classroom [Cag19, CR08, vB10]. Claude [Law16]. Clavius [Par17]. Claypool [McQ20]. Clements [Den15]. Clerk [Flo13, McC11, Ric14, Sim14]. Clifford [Bus18]. Codes [Ano19d]. Cofman [Sch22]. Cohen [Hit21]. cold [Ran21, Led16a]. Colin [Bru20, Wes20]. collaborative [Rei23]. Collection [Ano23l]. College [CK07, Flo16, McC16, Ric15a, Flo16]. collinearity [RB16a, RB16b]. Collins [Bee17]. Colloquium [Ano19e]. colonial [Mac20, WT12]. columns [Zel19]. combinatorial [Wil08]. combinatorics [Wil14]. common [Led19]. Communication [Led16a]. Commutativity [RB16a, RB16b]. compendii [Cra15]. complex [Lew12]. compounding [Bel15b, Big13]. Computation [Daw18, Ano23i, FR14]. computer [Bow15, SW17]. conceptual [Ano23o]. concerning [Pri08]. conference [PR18, Ste06, Whi15]. Congress [Emi12]. conjectures [Bur13]. Connecting [Cla10]. construction [Ano23e, PCMA15]. constructions [Lor12]. contemporaries [Wes20]. contemporary [Mal08, Man12]. contents [Rod17a]. continued [Gra09]. Continuity [BBG^{+}20, Har16]. continuous [Big13]. contribution [And09]. contributors [Zel19]. contrivance [Bra15].

Cooke [Hol18b]. copy [Mok18]. Cork [Hit21]. Correspondence [Poo15]. Cosmos [Car18]. count [Cla20]. Counting [Fro15, Per15]. counts [Wil14]. course [Dew17, dB09]. Craik [Fal20b]. created [Whi14]. creating [Dav17]. creation [And09]. crisis [Fri22, Kid22]. critical [Ano23f]. Cryptography [Ano19d, Wes16, Jan08]. CSHPM [Wil15g]. Cuisenaire [Cag19]. Cullen [Dau20]. culture [Hee19, PO13, Wes12]. Cunha [Dom14]. Cunningham [Bus18]. curious [Cra16]. currency [Big09]. curriculum [Krü10]. curve [Ano23e, Led19, SK18]. curves [Ber21]. cyclic [Atz15]. cyphering [Den15].

D [Fal20b]. Dan [Big19]. Dana [Cra16]. Danesi [Cri20a]. Daniel [Tak17, MS22]. Danilo [Spe19]. Danish [PR18]. D’Arcy [Hin24]. darkest [Son17]. Darwin [Led19, Ste09]. Dasgupta [Bow15]. David [Fea18, Sia15, Ano23j, DB21, Fin19]. Davis [Ald21]. Day [Wil15c, Cri15]. days [Big19]. deaths [Man15b]. Debate [Sta15]. DeBevoise [Law16]. December [Ano23j, Cri16, Ric15b]. decline [AH08]. Decoding [Big21, HS15]. Dedekind [Haf19]. Deduction [Daw18]. deductive [Pet15]. Dee [Law11, Ram11a]. Defining [Har16]. definition [Hol14, Nak14]. deification [Bel14]. Dello [Mar06]. delusions [Mur16]. Demidov [Hol18b]. demonstrativa [Cra17]. Denisova [McQ20]. derivation [Rec21]. derivative [PCMA15]. Descartes [Dea11]. Descriptio [Sim14]. description [Per15]. Descriptive [Gra20]. design [Krü10]. Desmond [Hit21]. Details [Pis16, PB16]. detective [Ams08, Gra09]. development [Ano23f, BK22, DB21, Par06]. devices [TCP19]. Devlin [Sim18]. Diaconis [Cri18a]. diagonal [Bra15]. diagram [Edw06]. diagrammatic [Fri19]. Diagrams [vM06]. Dicuil [RK19]. Did [Mok18, Nak14, ACO19]. Dido [Ach23]. differential [Ber21, Nak14]. Digital [Daw18, Ast20]. Dirac [CH16, Cri18b]. Dirichlet. [Sch21]. Discovery [Bus18]. discrete [Ast20]. discussion [Row15]. discussion-led [Row15]. Dispelling [Gra07]. diverse [Seg20]. documents [CO11]. Dodgson [Pec20, Wil18b]. Domenico [Rob18]. done [Man12]. drawings [Huy15]. drove [AC13a]. drums [TCP19]. Dualgruppen [Haf19]. d’Ulm [Fin19]. duplication [CC13]. during [Bai21, Led16a, Per15]. dynamic [Cag16].

e-book [Sch21]. E. [Hol16a]. Earle [Man16b]. Early [Ano19a, Ano20d, Bur16, Ano23a, Bee20, Bel15a, Bur13, Cal18, Cra17, Dau20, DB21, Hee19, Hig19, HMR17, Jam11, Jen10, Kar07, KJB11, Krü10, MG11, Per15, Ric16, Sat21, Tou19, War14, YO17, Ano23l]. earth [KJB11, Rec21]. Eclipse [Ken22]. ed [Sch15]. Eddington [Ken22]. Edinburgh [Har09]. edited [Law18, Nas16, Poo15, Sch15]. editing [Rob15]. edition [PB16]. Éditions [Fin19]. Editor [Bur15, Dru16]. Editorial [Ano06d, Ano14c, Ano15, Ano17a, Ano23g, Fal20a, Fal22, Fal24, Har21, Man13, Man14a, Man14b, Man15b, Man15d, Man15c, Man16a, Man16b, Man17, Ste06, Ste08, Ste10a, Ste10b, Ste10c, Ste13, War18a, War18b, War22]. editors [Man15a]. Edmund [MM19]. eds [Hig19, Hol18a, Ric14]. Education [PR18, Ano23f, ADD16, Bal15, Bar09, Dav18, Har09, HMR17, Kar07, Kas07, WT12, YO17]. educational [Row15]. Educator [BG10]. Edward [HS15, SK18]. effectiveness [Jan14]. Egyptian [SS22]. eighteenth [Dec11, Rei23, Big22]. eighteenth-century [Dec11, Big22]. Einstein [Ken22]. Ekaterina [McQ20]. electrons [Ran21]. elegant [Wes24]. Elementary [Cla10]. Elements [Sia15, Ano23a, BG06, Bel15b, Ram11a]. Elena [McQ20]. eleventh [Edw06]. eleventh-century [Edw06]. elimination [Ano23o]. Eliot [Bal15, Lee16]. Elizabethan [Ram11a]. Elizabethanne [Hig19]. Ellerton [Den15]. Elli [Rei23]. Ellis [BRO24, Ver14, Ver16, Ver18]. Elwes [Cra16]. Emerging [Ano17b, Man17]. Emerson [Gow06a]. Emil [Sch15]. Émilie [Opi16]. emphasis [MS22]. empiricist [Ver14]. encounter [Pri08]. encyclopedia [Cra06]. end [Big09]. Engine [Pad17]. engineers [Bar09]. England [Bee17, Jam11, Jen10, War07]. English [BG06, Bel15a, Hil14]. Enlightenment [Bai21, Led16b, Rod18, Wil18a]. environment [Cag16]. episode [Rau16]. episodes [Kar18]. epistemic [Gul22]. epistemology [Spe19]. equation [Woo19]. equations [Cra16]. Erdos [Rek19]. Essay [Ano23m, Ano23b]. essays [Led19]. Euclid [Rod18, Sia15, Ano23a, BG06, Bel15b, Ram11a]. Euler [McQ20, Wil18a, Atz15, Bar09, BG10, BWH22, Coa08, Gra08, Tou19, Wil08, Wil14, Woo19]. Eulerian [Mal08]. EUR [Fin19]. Europe [Ano23l, Gul22, Not20, Par17, War17, Abr18, Hig19]. Eustis [Cla10]. evaluating [SGPBGL22]. event [Tur15]. events [Ano06h, Ano06i]. ever [KS17]. everything [Ram11a]. evidence [Bur13]. evolutionary [Led19]. exact [Fri22]. excerpts [ME17]. exchange [Big09]. exchequer [CC13]. exclusion [MS22]. expectations [Rob15]. experimental [Wes24]. Exploring [Cag16, Cuo12]. exprimenda [Cra17].

F [Law18, Ver16]. facile [Cra17]. fair

[Sat21]. family [Led19]. famous [EMSS22]. fascism [Hol18a]. Father [CC13]. Fatou [HG12]. February [Ano23j, McC16]. Feingold [Hig19]. Feke [Van19]. Fenny [Man15a]. Fermat [Ano23e]. Ferreirós [Wag18]. Festival [Wil15a]. Fibonacci [Sim18, Par17]. fiction [Bel10, Gau10, Man10, Man12]. fifty [WF21]. figurate [Bee09]. figures [vM06]. Filling [vB10]. Film [Wil15c]. finding [Son17, Sim18]. findings [Mac20]. First [Bus18, Cam23, Emi12, KS17, Wil17a, WF21, YO17]. fistulas [Cra17]. Flemish [dB09]. flight [Hin24]. Flood [Ric14]. Flora [Har09]. Florence [Mag12]. Florentia [Tha10]. Florida [CH16]. flow [MG11]. fluid [Tou19]. fluxion [Ber21]. focus [Sel08]. Folding [Sha16, Fri16]. Folkton [TCP19]. footsteps [Wil14]. forgotten [Bra15, Rod17a, Sim18]. Formal [Daw18]. formalization [Fri19]. formula [Atz15, SS22]. Formulating [Bee09]. Formulations [MS22]. forward [AC13a]. found [Ano23o]. Foundation [Sim14]. Fountoukli [Tha10]. frameworks [BBE^{+}21]. France [Hol18a, KS15]. Francine [Law18]. Francis [Ano19a, Ano20d, Gow06a, Led19, Pri08]. Francisco [NF23]. Frankfurt [Sat21]. Frédéric [Hol18a]. Freguglia [Cal18]. frequentism [Ver14]. Frigyes [Rod17b]. fruit [Bal15]. function [PCMA15]. functions [Coa08, Har16]. funding [Row15]. Funicularia} [Ala20]. future [Nas16, Nas16].

G [Lew12, Lew16]. Gabriella [Rob18]. Galilean [Spe19]. Galileo [PO13]. Gallery [McC18]. Galton [Ano07d, Led19, Pri06, Pri08]. Game [Big18, CRPH19]. games [Sil11]. Gareth [Man15a]. Gaspard [Gra20]. Gaussian [Ano23o]. Geddy [Cra15]. Gems [Ote21]. Generative [Row24]. genesis [Bow15, Bra13, Haf19, Not20]. Geneva [PB16]. genius [Sim18, Wil18a, vWW18]. Geometric [Bur07, Rod18]. geometrical [Lor12]. géométrie [Dea11]. geometry [Ano23a, Bee20, Cag16, Cra20b, Fri16, Gra20, Gul22, Tay11, Wil17a, Wil17b].

George [Hit21, Bal15, Cra20a, Gow06b, Led19, Lee16, Mok18, Mur16, Sim08, Ste09]. German [McQ20]. Giaquinta [Cal18]. Giovanni [Rob18]. Giraldes [NF23]. glass [Ano23o]. Glen [KJB11]. gnomons [NT16]. go [MM19]. Goals [Lee13]. goes [Bar09]. Goldbach [Rod17a]. Gongard [Law16]. Good [Sch15, Law06]. gossiped [vWW18]. got [Har09]. Göttingen [Hol20]. Gould [Ken12]. Grande [Fin19]. Graph [Cam23, Big18, Mag12]. graphics [Mag12]. Graphs [Ano23h, Ano24a]. Grattan [Ric15b, Man15b, Man15d]. Grattan-Guinness [Ric15b, Man15b, Man15d]. Great [Daw18, Bee20, Cri18a, DB21, HG12]. greatest [War17]. Greek [Cuo12, Kas07]. Gregory [Ams08, Cri20b, DB21]. Gresham [CK07, Flo13, Flo16, Ric15a, Flo16, Wil17a, Wil17b]. group [Fri19, Rau16]. groups [Sør13]. Guerre [Fin19]. Guinness [Ric15b, Man15b, Man15d]. Guy [Sta15].

H [And09, Ver16]. Hamilton

[Cra20b, Sim08, vWW18]. handsome [KS17]. Hans [GG14b]. Hard [Bay07]. hardback [Sch21]. Harriot [Mac21, Big13, GSttmoJS17, Sel08]. hb [McQ20]. health [Jen10]. Heavenly [Dau20]. heavens [KJB11]. Heesch [Rei23]. Heinrich [Rei23]. held [Emi12]. Henri [Mun12]. Henry [AC13a, Cra20b, Jag06]. her [Ano17b, Har09, Man17]. heritage [Mun12]. Hermann [ACO19, Fri16]. Hertfordshire [AC13b]. Heuraet [Ano23e]. higher [Har09]. Hilbert [Rei23]. Hill [Ste09]. Hippocrates [Cag16]. historical [BGI11, BGI14, Dav17, Gra07, Gui18, Jan14, RB16a, RB16b, SGPBGL22]. historically [Ano23o]. histories [Cra16]. histories/A [Cra16]. historiography [vB10]. History [Ano17b, Ano19d, Ano19e, Ano24b, CC13, Hol16b, KT10, Man15c, PR18, Ric15a, Sim14, Sta15, Whi15, Wil15c, Ano07k, ADD16, Cag19, Cla10, Cra16, Den15, EMSS22, Fad15, Hol06, Jan08, Kay06, KJB11, Law06, Law11, Man12, Man17, McD18, McQ20, Neu08, Ric16, Rog20, Row24, Son14, Sør17, Ste19, Wil15d]. Holliday [Gow06a]. Holmboe [Chr09]. Hong [WT12]. Honoring [Cla20]. honour [dB09]. Hot [Ran21]. House [Led16b]. humour [War07]. hundred [Cam23, Wil17a, Wil17b]. Hunger [Ric16]. hunt [Min09]. Hutton [War17]. hydraulics [Cra17].

Ian [Cra16]. Iceland [Sch15]. iconography [Mag12]. idealist [Ver14]. ideas [Cri18a, ME17, RB16a, RB16b]. identities [Hol06]. II [Nas16, RB16b]. ill [Law16]. Ilona [Hol15]. IMA [Wil15e]. Images [Hol18a]. imagination [Rod18]. impact [Row15]. imperial [Dau20]. implications [Row15]. Impressions [Ken22]. Impressões [Ken22]. inalienable [Abr20]. inclusion [MS22]. incommensurabilities [NT16]. inconvenience [Wes24]. India [Agg07]. individual [Big19]. industry [Bel15a]. Inference [Law18]. infiniment [Bra13]. Infinite [Ano23i, Sia15, MM19]. infinitesimal [BBE^{+}21]. infinity [Rau13]. Influence [Ano17b, Man17, Smi08]. influences [Cra20a, Tou19]. informal [CH16]. informed [Ano23o]. Insight [Haf19]. inspiration [Pri06]. Institute [Led16a, Wes16, Tur15]. Instrument [SGPBGL22, Cra10]. instruments [Kid22]. insurance [Bel15a]. integrable [PCMA15]. integral [Ano23e]. Intellectual [Sta15, Mun12]. intended [Ano23a]. interconnection [RB16a, RB16b]. interconnections [BK22]. International [Prz09, Emi12]. interplay [Pis16, Wag18]. interpretation [Gui18, Ver14]. introduced [Ric19]. Introducing [Ber21]. introduction [Lew12, Per15, Son14]. intuition [Fri22]. invariants [Par06]. inventioni [Seg20].

Investigating [Hol14]. investigation [Dea11, Dew17]. investigations [CH16]. Ireland [Flo06]. Irish [Whi15]. Isaac [Bee20, War10]. ISBN [Cla20, Cri20a, Dau20, Fin19, Hit21, Mac21, McQ20, Ran21, Sch21]. Ishango [Mac20]. Isles [Ald21, Dav18]. issue [Law18]. issues [BBG^{+}20]. Italian [Hol18a, Cri20b]. Italy [PO13]. Italy/Galileo [PO13]. IV [Poo15]. I’ve [KS17]. Ivor [Man15b, Man15d, Ric15b].

J [Nas16, Poo15, Ric16, Sen16]. Jackie [Ano17b, Man15b, Man15d, Rob15]. Jacob [Ala20]. Jacqueline [Son14, Van19]. Jakie [Man17]. James [Cri20b, Flo13, McC11, Sim14, Ric14]. Jan [Daw18]. jar [Man10]. Jean [Sta15]. Jean-Guy [Sta15]. Jean-Pierre [Sta15]. Jesuits [dB09]. Jitsuro [Tak17, Tis13]. Joel [Cra16, Fai17]. John [Ano24a, Poo15, Whi14, AH08, Ano23h, Bee17, Cra15, DB21, Gow06a, Hil14, Man16b, Pri08, Ram11a, Sim14, Ver14, Wes16]. Johnny [McD18]. Johnston [Sen16]. Joint [Man15c, Sim14, Wil15e, Wil15g, Ric15a, Tur15]. Jones [Cla20, Car18]. José [Wag18, Dom14]. Journal [Sch15, Law18]. journals [Prz09]. Journey [Fro15]. Juan [CRPH19]. Judita [Sch22]. Julia [HG12]. July [Led16a]. June [Led16b, Ric15b, Wes16].

Kaliningrad [Mal08]. Kant [Ver16]. Kaplan [SK18]. Karen [Ric16]. Karin [Sch15]. Katya [McQ20]. Katz [Ric16]. Kazimirovich [Hol09]. Keith [Sim18]. Kelvin [Cra12, Row14]. Ken [Den15]. Kepler [Gul22, Rec21]. key [Cra16, Jan08]. Kim [Nas16]. King [Cra15, Sia15]. Knight [Cra13]. knowledge [Bal15, Bee17, Wag18]. Kong [WT12]. Kovalevskaya [Ano23n, KS17].

L [Pec20, SK18, Wil18b]. laboratory [BRO24]. lady [KS17]. Largs [Cra12]. late [Bee20, PO13]. later [Bur13]. Latin [Hol18a, Ric19]. latitude [Son17]. Lattice [Boa07]. lattices [Haf19]. Lavant [TCP19]. Laws [Mok18]. learning [Fad15, Wes12, NK15]. Lecture [Bur06, Flo13, CK07]. lectures [Cra12, Ste09]. led [Row15]. legacy [Gra20, Tak17, Tis13, Cri20a]. Legendre [MS22]. Leibniz [ME17]. Leibnizian [Jes15, BBE^{+}21]. L’Élite [Fin19]. length [Ano23e]. Leonardo [Huy15, Pis16]. Leonhard [Bar09, BWH22, Tou19, McQ20, Wil18a]. Leslie [Ver14, Ver16, Ver18]. Lessons [Krü10, Law11]. Let [War10]. Letter [Bur15, Dru16, Ano07d, Ver18]. Letters [Gow06a, McQ20]. Lëvshin [Hol18b]. Levy [Cra16]. Levy [?]TheCrampin:2016:NSR. Lewis [Law18, Pec20, Wil18b, Law18, Mok18, Wil18b]. l’Hôpital [Bra13]. liberated [Abr20]. Life [Gow06b, Man15a, Bel15a, Gra08, Kar18, Mac21, McC11, Ric14, Van19]. light [Hit21]. limit [Nak14]. limit-avoiding [Nak14]. limits [Bur06]. line [Ams08]. Linear [BK22]. lines [vM06]. Lisbonenses [NF23]. Lister [Ano24a, Ano23h]. literature [ADD16, Rod18]. lives [Sør17]. LMS [Cri15]. Local [CC13, Cla10, Wes20]. location [BGI14]. Logan [Ver16]. Logarithmorum [Sim14]. logarithms [Ano23i, Bur16, FR14, Whi14]. logic [Opi16]. London [Flo16, Fai17, Sat21]. long [Rod18]. look [FR14, SS22, vW22]. looking [Ano23o]. looks [Kun07]. Lord [Row14]. lost [Min09]. Loterias [NF23]. Louis [Fai17]. Lovelace [Flo16, HMR17, Pad17]. Luca [Huy15, Pis16, Smi08]. Ludlam [Dec11]. lunar [Ste09]. lunes [Cag16]. Luzin [Hol18b].

M [CH16, Cla20, Den15, Hol21, And09].

MacHale [Hit21]. machine [Ast20]. Machinery [Daw18]. Mackenzie [Cra16]. Mackenzie [?]17Crampin:2016:NSR. Maclaurin [Bru20, Wes20]. MacMahon [Wil14]. MacTutor [Hol17]. mad [vW22]. main [RVME18]. Making [Mar06, Hol16a, MM19]. Malcolm [Law16]. Malfatti [Lor12]. man [MM19]. Manchester [And09]. manifest [McC18]. manipulatives [Cag19]. Manucci [Smi08]. manuscripts [GSttmoJS17]. Marcel [Cri20a]. March [Ano17c]. Mariano [Cal18]. Mark [Ric14, Kar18]. marriage [Rau16]. Martin [Ano24a, Ano23h]. Mary [Opi16]. material [Wes12]. math [Kle07]. mathematic [Row15]. Mathematical [Agg07, Ald21, Ano19e, Ano23l, Cla20, Dav17, Fea18, Fri19, Kid22, Led16a, Mus10, Sat21, Wes16, Wil18b, Ano23b, Bar09, Bas20, Bee17, Bla08, CH16, Cra10, Cra16, Dav18, EMSS22, Fri22, Gui18, Hil14, HG12, Hol16a, HMR17, Jes15, Kas07, KS17, Law16, Mag06, Mar06, ME17, Pad17, Pec20, PO13, Pet15, Pis16, Prz09, RB16a, RB16b, Sch21, Sil11, Sim18, War07, Wil18a, vM06, Hol17, Led16a, McD18, Wag18]. Mathematicall [Ram11a]. mathematici [Cra15]. Mathematician [Dav18, AC13a, Big22, DeC16, Dec11, Ken12, Man15a]. Mathematicians [Bel15a, Emi12, Tur15, And09, Gra07, KS15, Law06, NK15, Ric19, War17, Cla20]. mathematicisation [Bai21]. Mathematics [Ano17b, Ano18a, Ano19f, Ano23l, Bar09, Bel10, Big09, Flo06, Flo16, Fro15, Hol15, Hol18a, Jen10, Led16b, McC18, PR18, Whi15, Wil15c, Wil15f, WT12, Abr18, Abr20, AH08, Ano07k, Ano23f, ADD16, Bal15, Bra15, Cag19, Chr09, CR08, Cla10, Den15, Dew17, Fad15, Gau10, Gra06, Hee19, Hol20, Jan14, Kar07, Kay06, KJB11, Krü10, Law06, Law11, Lee16, Mac20, Man10, Man12, Man17, Mon11, MG11, MF22, Nas16, PO13, Ram11a, Rei23, Rek19, Ric06, Rog20, Row24, SGPBGL22, Sch22, Son14, Sør17, Ste19, Van19, War14, Wes12, Wes22, Wes24, Wil08, Wil14, Wil15d, YO17, Zel19, dB09, vB10, Ano24b, Man15c, Sim14, Ste06, Wes16, Man17, Cra16, Woo19, McC18]. mathématiques [Fin19]. Maths [Law06, Gra07, Cra16]. matter [Gra07]. Matthew [Rod18]. Maxwell [Flo13, Ric14, Sim14, McC11]. May [Hol16b]. Maya [Mon11]. Mayor [RVME18]. McCartney [Ric14]. me [Man12]. meandered [Kar07]. Measuring [TCP19]. mechanics [Seg20, Tou19]. Mechanism [Car18]. medieval [Oak22, Per15, Not20]. Mee [Cra06]. Meeting [Ano19b, Ano20a, Ano20b, Ano21a, Ano21b, Ano22a, Ano22b, Ano23d, Man15c, Sim14, Ste23, Wil15b, Ano18a, Ano22c, Ano23c, Ano24c, Cri16, Flo16, Ric15a, Wil15e, Ano19c, Wil15g]. Meetings [Ano09b, Ano10a, Ano10b, Ano11a, Ram11b, SPR09, War09, War11, Wee10, Ano06a, Ano06b, Ano06c, Ano06g, Ano06h, Ano06i, Ano07a, Ano07b, Ano07c, Ano07h, Ano07i, Ano07j, Ano08a, Ano08b, Ano08c, Ano08g, Ano09a, Ano10f, Ano10g, Ano11b, Ano12a, Ano12b, Ano12d, Ano13a, Ano13b, Ano13d, Ano14a, Ano14b, Ano14d, Fal18, SW12, Wes18, Wil13]. Meier [SK18]. memoirs [Ano23n]. memories [Rob15]. men [BG06]. Mengoli [ME17]. mental [Jen10]. meridian [Ams08]. Merry [War07]. Merzbach [Sch21]. Mesopotamian [Man23]. metabiographical [Sør17]. Metaphorical [Fri22]. metaphors [Ano23o]. Methodi [Cra15]. Methods [Sen16, Ano23b]. Michael [Chr09, Nas16]. Michèle [HG12]. mid [KS15]. mid-twentieth [KS15]. middle [Law11]. Mielke [Sen16]. milestone [McQ20]. military [Bar09]. Mirifici [Sim14]. misconception [Son17]. Mistakes [Pri08]. mitraille [Fin19]. models [Lew16]. Modern [Ano23l, Rob18, BBE^{+}21, CR08, Fre13, Hee19, Hig19, Jam11, Not20, Per15, Sat21, Spe19, War14]. module [Jan08, Jan14]. modules [Haf19]. Moktefi [Law18]. molecules [Ran21]. Molyneux [Ano24a, Ano23h]. Monge [Gra20]. monochord [Bel15b]. Montel [HG12]. Mordechai [Hig19]. Mordell [Fai17]. Morgan [Cri15, McQ20, Big22]. Moscow [Hol15]. most [Woo19, vWW18]. motivations [Ano23b]. MR3169973 [GG14a]. Much [BG06]. multiplication [Boa07]. multivariate [BK22]. muse [PO13]. Museum [McC18]. Music [Wes16, Ano23a]. musings [vB10]. Mutio [Mar06, PO13]. Mysticism [Hol15]. myth [Hol16a, Not20]. myth-making [Hol16a]. myths [Gra07].

N [McQ20]. Nagura [Tak17, Tis13]. Nahin [Ran21]. Napier [FR14, Sim14, Whi14]. natural [Bur16]. necessary [BG06]. neglected [Cra20a]. Neil [Bur06]. Nepal [Ano23f, Bas20]. Nerida [Den15]. networks [Big17]. Neumann [Ano24d, Hol21]. never [Rau13]. Newman [And09, NB21]. News [Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08d, Ano08e, Ano08f, Ano09c, Ano09d, Ano10c, Ano10d, Ano10e, Ano12c, Ano13c, Ano17c, Ano19b, Ano20a, Ano20b, Ano21a, Ano21b, Ano21d, Ano22a, Ano22b, AFGM09, FC20, Led21, Mar14, Ste14, Ste23, Ano22c, Ano23c, Ano23d, Ano24c]. Newton [Hig19, Bee20, Bel14, PB16, War10]. Newtonian [Opi16, Bru20, PB16]. next [Wil17b]. night [Son17]. Nightingale [Mag12]. Nikolaevich [Hol18b]. Nikolai [Hol18b]. nineteenth [Agg07, Ano23b, Dom14, Gag17, Gra06, Jen10, Kar07, Lor12, Zel19, Ano18a]. nineteenth-century [Gag17, Jen10, Lor12]. Non [Wil15f, PCMA15]. non-Riemann [PCMA15]. Non-Western [Wil15f]. nonconstant [PCMA15]. normal [Led19]. normaliens [Fin19]. North [Den15, Wil15e]. North-West [Wil15e].

Note [Boy24, Cra16, Dav18, Rau13, Man15b]. notes [Kar07, PB16]. nothing [Oak22]. Notices [Ano09c, Ano06e, Ano06f, Ano07e, Ano07f, Ano07g, Ano08d, Ano08e, Ano08f, Ano09d, Ano10c, Ano10d, Ano10e, Ano12c, Ano13c, AFGM09, Mar14, Ste14]. Nourse [Gow06a]. Nova [Rec21]. novels [Bal15]. Novembertagung [Ste19]. Number [Ano05a, Ano05b, Ano05c, Hol16b]. numbered [Big19]. numbers [Bee09, Cag19, Dau20, Dej14, Jam11, Lew12, McD18, RK19, Sta15]. numeracy [Cuo12, Kay06].

o [Ken22]. o/and [Ken22]. OBE [Hol21]. Obituary [Ano23j, Ric15b]. Observations [Huy15]. October [Flo16]. Oddi [Mar06, PO13]. Old [Fre13, Dej14, Llo12]. one [War17, Wil17a]. opening [Dav18]. orbit [Rec21]. ordinary [Zel19]. Origins [Daw18, Jes15]. Ostwald [Nas16, Nas16]. Ostwald/Architecture [Nas16]. Other [Ano06g, Ano06h, Ano06i, Ano07h, Ano07i, Ano07j, Ano08g, Ano10f, Ano10g, Ano12d, Ano13d, Ano14d, War09, War11]. Oulipo [Bel10]. our [Big19, Emi12, Kar07]. Overstating [Gra06]. own [Mok18]. Oxford [Cri20a, Dau20, Led16a, Led16b, Mac21, McC16, Gag17, Wes16].

P [Sen16, Lew12, Lew16]. Pacioli [Huy15, Pis16, Smi08]. Palimpsest [Lee13]. pamphlet [Cra17]. Paolo [Cal18]. paper [Tak17]. papers [CH16, Hil14]. Paradox [Law18]. Parr [Cra20b]. Parshall [Ric16]. Part [Cra13, RB16a, RB16b, Wil17a, Wil17b]. partial [Man10]. participation [MF22]. partisanship [Kle07]. party [vW22]. pass [ACO19]. Passedix [CRPH19]. passing [Man15d]. past [Sør17]. path [Spe19]. paths [Cri20b]. patron [NK15]. Patronage [Ano19f]. Patterson [DeC16]. Paul [CH16, Cri18b, Rek19]. pb [McQ20]. pedagogical [Dea11]. Pell [Hil14]. pentagon [Sha16]. Perceptions [Ano07k]. period [Cal18, Hee19, Per15]. Permutation [Sen16]. Persi [Cri18a]. perspective [ADD16, BGI11, BGI14]. Perspectives [Ric14]. pestalozzianism [Kas07]. Peter [Hol21]. petits [Bra13]. Philip [Poo15]. Philosophy [Man15c, Bai21, Rei23, Van19]. Physicists [Tur15]. physics [McQ20, Tur15]. picture [Hol20]. Picturing [Pad17]. Pierre [Sta15]. pilot [Cla10]. pioneering [Woo19]. Plath [Man10]. Plato [Daw18]. Platonic [Llo12]. Playfair [AH08]. playfulness [Gul22]. Plot [Ano24a, Ano23h]. plumbeas [Cra17]. plurality [Par17]. pneumatics [Cra17]. poetic [McC11]. poetry [Bas20, Bel10]. Poincaré [MS22, Mun12]. point [Fad15]. points [vM06]. Poland [Prz09]. polar [Mag12]. Political [Sta15, Kle07, YO17]. Polyakhova [McQ20]. polygonal [Cag19]. polytechnic [Gra20]. Poor [War07]. popular [Cra12]. Portable [Car18]. portrait [Dec11, HS15]. portrayals [KS17]. positive [BRO24]. post [Spe19]. post-Galilean [Spe19]. postulating [Abr20]. pp [Cri20a, Fin19, Mac21, McQ20, Ran21, Sch21]. Practical [Fro15]. Practice [Ano18a, Bru20]. practices [Wag18]. praeface [Ram11a]. praxis [Sel08]. Pre [Fad15]. Pre-service [Fad15]. prediction [Row14]. prehistoric [Man12]. prerogative [Abr20]. present [Bra15]. Presente [Gui18]. Press [Cri20a, Dau20, Hit21, Mac21, Ran21]. Prévost [Sta15]. primary [Law11]. Princess [McQ20]. Princeton [Ran21]. Principia [PB16]. principle [MS22]. Printing [War14]. priority [BBG^{+}20, Hol14]. Pritchard [Ano07d]. Prize [Ano23m, Ano24d, Ano23b, Ano19a, Ano20d, HG12]. Prizes [Ano19g, Ano20c, Ano21c, Ano22d, Ano23k, Ano24b]. probability [DB21, Ver18]. problem [Ach23, Ala20, Kun07, Lor12, Rei23]. problems [BGI11, BGI14, EMSS22, Mas07, Pet15, YO17]. Procedures [BBE^{+}21, Ano23i]. product [Ber21]. Production [Ano23l]. Professor [Fal20b]. Professors [Wil17a, Wil17b]. Progress [Ano18b, McC16, Smi15, Wal14]. progression [Den06]. progressions [Bur07]. project [Cla10]. promoting [Gra07]. promotion [Bee17]. proof [Boy24, BRO24, Tan20]. proofs [NT16, PB16]. proposal [Ano23a, Ano23f]. Propositions [Bel15b]. prosopographical [Dav18]. Protection [NK15]. Ptolemy [Min09, Van19]. Public [Sta15, Jan08]. public-key [Jan08]. publication [Sim14]. Publications [Ano09f, Ano06j, Ano06k, Ano06l, Ano07l, Ano07m, Ano07n, Ano08h, Ano08i, Ano08j, Ano09e, Ano09g, Ano10h]. publike [Bee17]. published [Prz09]. publisher [Gow06a]. put [War10]. puzzle [Bja09, Zel19]. pyramid [SS22]. Pythagora [Cri20a]. Pythagoreans [Hol15].

quadrature [Ber21]. quadrilateral [Atz15]. Quantifying [Bai21]. quarter [Kle07]. quaternions [vW22]. Queen [McC16]. Quesiti [Seg20]. quest [Sim18]. Quincunx [Ano07d, Pri06].

radical [Big22]. rank [SW17]. Raphael [PO13]. rare [Cra10]. ratios [Bel15b]. Raymond [Ric14]. reactions [Cra20a, Fri22]. readers [Cri20b]. Reading [Ano23l, Jam11, Hig19]. real [Bur06, Har16, YO17]. really [SW17]. reason [Tay11]. reasoning [Fri19, Pet15, vM06]. reconstruction [Fri22]. Recorde [Man15a]. records [Big21]. recreational [Zel19]. rectangles [Man23]. rediscover [Sim18]. rediscovered [Rod17a]. reflections [YO17, Gra06]. reformers [Ano23b]. refugee [NK15]. regular [Sha16]. reinterpreting [Spe19]. remark [Tak17].

Remarks [Ver14]. Renaissance [Ano18a, PO13]. rendition [Ano23f]. repercussion [Dom14]. reports [Ano19c]. Research [Ano18b, Ano19a, McC16, Smi15, Wal14, War14, ADD16, Rei23, Row15, Sch22]. resembling [Bur16]. resource [Fre13]. resources [Dav17, Law06, SGPBGL22]. response [Ano07d]. Restoration [Bee17, War07]. Rev [Ver16]. Revealing [Car18]. Review [Abr18, Big19, Bow15, Bus18, Cal18, Car18, Cla20, Cri18a, Cri20a, Dau20, Daw18, Den15, Fea18, Fin19, Hig19, Hit21, Hol15, Hol18a, Hol18b, Jes15, KJB11, Law16, Mac21, Man15a, McD18, McQ20, Nas16, Opi16, Pec20, Poo15, Ran21, Ric14, Ric16, Rob18, Rod18, Rog20, Sch21, Sch15, Sen16, Sia15, Sim18, Son14, Spe19, Sta15, Tan20, Van19, Wag18, Wil18a, Woo19, McC18]. Reviews [Ano06m, Ano07o, Ano07p, Ano07q, Ano08k, Ano08l, Ano13e, Ano14e, Ano14f, CMBV12, GDR^{+}11, MCC^{+}09, MGW^{+}10, MH11, MMP^{+}09, OHL12, PHN^{+}09, SR13, TMIvM10, WM10]. revisited [Cra06, FR14, Per15]. revolution [Jes15, Opi16]. revolutionary [DeC16]. Rewley [Led16b]. Rewriting [Den15]. rhyme [Bja09]. Richard [AC13b, Cra16, Jes15]. richer [Hol20]. Riemann [PCMA15]. Riesz [Rod17b]. rinded [Bal15]. ring [Hol14]. rise [Bee20, Big19]. risk [Big19]. Risorgimento [Hol18a]. Robert [Ano24a, AC13b, Ano23h, DeC16, Sat21, Ver14, Ver16, Ver18, Man15a]. Roberts [Man15a]. Robin [War07, Woo19]. Robins [Bar09, Wes24]. Robyn [Mac21, Opi16]. rods [Cag19, Mac20]. Roger [Hol18b]. role [Bai21, Den15, KS17, MG11, Ver16]. roles [DB21]. Roman [Cuo12]. Ronald [McQ20, Wil18a]. Root [Bur13, Man23]. Routing [BGI11]. Row [Fri16]. Rowan [Sim08, vWW18]. Royal [Hil14]. RSA [Jan08]. Rue [Fin19]. rule [WT12]. Russell [Gau10]. Russian [Hol15, Hol18b, Hol14, Kar07].

S [Hol18b, McQ20, Wil18a]. Said [Law18]. sale [Mok18]. Salmon [Gow06b]. sample [Pet15]. Sanskrit [Pet15]. satanic [Not20]. scale [Bra15]. scholar [Kar18]. scholarly [Dew17]. School [Cla10, Den15, Law11]. schoolboy [Lew12]. schooling [Kar07]. Schools [Ano23m]. Science [McC18, NK15, Wil15a, Fri22, Seg20, Bow15, Mac21, Spe19]. sciences [HG12]. Scientific [Car18]. scientists [SW17]. scope [Lee13]. Scotland [MF22]. Scottish [Rod18, Bai21, Bee20, Cra17]. Scriba [GG14a, GG14b, Poo15]. Scrots [HS15]. search [CH16, Cra13, Ken12]. Seduced [Opi16]. seen [KS17]. Selections [Seg20]. Semitism [Hol15]. senses [Tay11]. sequence [Man23]. Serbs [Dej14]. Sergei [Hol18b]. service [Fad15]. set [Big18, Lor12]. seventeenth [Big17, Gul22, Krü10, dB09]. seventeenth-century [Big17, Gul22]. several [Kar18]. Sex [Man10]. shade [War10]. shall [Emi12]. Shandy} [Gau10]. Shannon [Jan14]. Sheet [Ano17c]. Shelly [Cla20]. short [Gra08, Son14, vB10]. should [Gra07]. Sieve [Tis13, Tak17]. Silva [MS22]. Simpson [Ach23]. Sinclair [Cra20a]. sine [Ano23n]. Singmaster [Ano23j]. Sir [MM19, vWW18]. sisters [Hol18a]. sive [Cra15]. sixteenth [RVME18, Seg20]. Sixth [Man15c, Wil15g]. Skyrms [Cri18a]. Slavs [Dej14]. Smith [Man15a]. Smyrna [Cag19]. snow [Gul22]. Social [Sta15, Rek19, Tha10, YO17]. Society [Ano24b, Cla20, Man15c, Sim14, Ano21d, Hil14, Led21, NK15, Wil18b]. Sofia [KS17]. solids [Llo12]. solve [Sch22]. solving [EMSS22]. Some [CO11, Cra20a, Cra16, EMSS22, vM06]. Somerville [Opi16]. sortes [BG06]. source [Rog20]. source-based [Rog20]. sources [RVME18]. sous [Fin19]. Spa [PR18]. Space [Sia15]. Spain [Ber21, RVME18]. Spanish [CRPH19]. special [Coa08, Law18]. spelling [NB21]. spirit [Abr20]. spread [Gra20]. squadro [Mar06]. square [Hol06, Man23, RK19]. squares [Ric19]. St [Ams08]. Stability [BRO24]. stage [War10]. stamps [Wil15d]. State [CH16, Sta15]. Statistical [Sen16, Big19, Lew16, Mag12]. Statistics [Ric15a, Bay07, BK22, KT10, Led19, Mag06, Sta15]. steam [Wes22]. Stedall [Man17, Ano17b, Man15b, Man15d, Son14]. Steve [Abr18]. Stewart [Cra16]. Stewart [?]MathsCrampin:2016:NSR. stone [Lor12]. Stonehenge [TCP19]. Story [Fea18, Ams08, Gra09]. student [KT10]. Studies [Ano23l]. study [Bra15, Fad15, Fri19, Jes15, RB16a, RB16b, Zel19]. Studying [Sør17]. style [Nak14]. Subrata [Bow15]. substitution [Ano23o]. Subverting [Rob15]. Sundara [Fri16]. Supplement [Ano04]. Surrey [Bla08]. Survival [SK18]. Suschkewitsch [Hol09]. Svetlikova [Hol15]. Sylvia [Man10]. Symbolism [Hol15]. Symbols [Wes22]. symmetry [Neu08]. symposium [War14].

T [Hol16a]. table [Per15]. tables [Bur16]. Tait [Lew12, Lew16]. tale [Hol16a, PB16]. Taming [Ric16]. tangency [Kun07]. tangled [Jes15]. Tartaglia [Seg20]. Tate [Sch15]. Taylor [Ano19a, MM19, Ano20d]. tea [vW22]. tea-party [vW22]. teacher [Fad15]. teachers [Kay06, Law06, Law11]. teaching [Fre13, Jan08, Jan14, Sch22]. techniques [Dea11]. Ten [Cri18a]. textbook [Mar06]. textbooks [Chr09, Den06, Fre13, McQ20, SGPBGL22, YO17]. texts [Bra15, Gui18, Pet15]. their [Coa08, DB21, Gra06, Law06, Zel19]. Theodoru [NT16]. Theon [Cag19]. theorem [Law16, NB21, Woo19]. theoretical [Pet15]. Theories [Daw18]. Theory [Cag19, Hol16b, Wes16, BRO24, Bru20, Cam23, Par06, Rau16, Ste09]. there [Ano23o, Bee20]. Thick [Bal15]. Thick-rinded [Bal15]. things [MM19, McD18, Wes22]. Third [Whi15]. thirteenth [Big09]. Thomas [Ach23, Big13, Cra13, GSttmoJS17, Mac21]. Thompson [Hin24]. Thomson [Cra12]. those [Bur16]. thought [Mok18]. three [BBE^{+}21, Kun07, Rod17a, Wil17b]. tide [Row14]. time [Fai17, TCP19]. timelines [Mus10]. Times [Man15a, Kid22, Bay07]. Tisdale [Tak17]. Tom [Sch15]. Tomàs [Ber21]. Tortoise [Law18]. tradition [Hol20]. training [Kay06]. trans [Law16, Sch15]. transcriptions [Lee13]. translated [Hol18b]. Travel [Ken22]. Treatises [Cra20a]. tree [Bal15]. triangular [RK19]. Tribute [Ano17b, Man17]. trigonometry [KJB11]. Tristram [Gau10]. truth [Gra07]. Tudor [Man15a]. Turkey [ADD16, Fad15]. Turkish [YO17]. twentieth [BK22, KS15, Ric16, YO17]. twenty [YO17]. twenty-first [YO17]. Two [Fri16, Rod17b, SS22, Zel19].

Undergraduate [Ano23m, Gag17, Ano23b]. Understanding [Cag19, Bur06, KT10]. unitis} [Emi12]. universal [Wes20]. universe [Cra16]. University [Cri20a, Dau20, Hit21, Mac21, Ran21, Wes16, Dav18, CH16]. unknown [CO11, Ric16]. unplanned [Row15]. Unquantifiable [Bai21]. unreasonable [Jan14]. untrodden [Cri20b]. USA [Dom14]. Use [Ano23l, Per15]. uses [Man10]. Using [Kay06, Ald21, Sør17]. Uta [Sch21].

V [Hol18b]. values [Coa08]. variables [Har16]. variance [KT10]. Variations [Tan20, Cal18]. Venn [Edw06, Pri08, Ver14]. Vernacular [Tay11]. version [CRPH19]. versions [Rod17a]. versus [Ver14, Wes24]. very [Son14]. VI [HS15]. Viagem [Ken22]. Viagem/Travel [Ken22]. Victor [Ric16]. Victorian [Flo06, Mag06, Mag12, Ste06]. views [Fad15]. VII [Seg20]. VIII [Seg20]. VIII.4 [Bel15b]. Villani [Law16]. Vinci [Pis16]. Viribus [Emi12]. Visual [Hee19]. vital [Mag06]. Vito [PCMA15]. Volterra [PCMA15]. Volume [Ano04, Ano05a, Ano05b, Ano05c, Nas16, Poo15, Rog20, SS22]. Vygodsky [Kar18].

W [Sen16]. walk [Bla08, Mal08]. Wallace [Cra10, CO11]. Wallis [Poo15, Wes16]. Walsh [Mur16]. Walton [Ver18]. war [Ken12, Kle07, Tak17, Tis13, Led16a]. Warner’ [Ano24a, Ano23h, Sel08]. wars [Prz09, Rod17b]. was [BWH22]. watchword [Emi12]. water [Wes20]. way [Van19]. weakness [Led19]. Web [CK07, Law06]. web-based [Law06]. Weierstrass [Nak14]. weights [Seg20]. Welwood [Cra17]. went [Rau13]. West [Wil15e]. Western [Ric19, Wil15f]. Weyl [ACO19]. Where [SW17]. Whewell [Ver16]. while [Gra07]. whist [AC13a]. Whitaker [Ric14]. Whitfeld [AC13a]. Whittaker [MM19]. Who [Ric19, SW17, AC13a, Cla20, MM19, Sim18]. Wickman [Rod18]. Wide [CK07]. Wiener [Fri16]. will [Jag06]. William [Ano24a, vWW18, AC13a, Ano23h, Big22, BRO24, Cra10, CO11, Cra12, Cra17, Dec11, Gow06a, HS15, Sim08, Ver16, Ver18]. Williams [Nas16, BRO24]. Wilson [Woo19]. wind [Wes20]. Winner [Ano20d, Ano24d, Ano23b]. Winners [Ano23m]. Winton [McC18]. without [vW22]. Woman [Dav18]. Women [Dew17, Flo16, KS15, MF22, Ald21, Cla20, Dav17, Dav18, Har09, Cla20]. Wonder [Car18]. Wonders [McD18]. words [Cra16]. work [Coa08, Gow06b, Ken12, Ric14]. Works [Daw18, Ano23e, Ano23i, Sat21].

Workshop [Ano19e]. World [Sch15, Wil18b, Ast20, CK07, Pec20, War14, YO17, CK07, Rod17b, Car18, Cra16, Sim18]. Wright [Son17]. wrote [Dej14]. Wussing [GG14a, GG14b].

X [Ano23a]. xiii [Cla20]. xix [Sch21]. xvii [McQ20].

years [BG06, Cam23, Wil17a, Wil17b, WF21]. young [Law06]. Yvonne [Hit21].

Zero [Oak22, Cra16]. Zeta [Coa08]. Zurich [Emi12].

References

Abrams:2018:BRA

[Abr18]  Ellen Abrams.  Book re-

view: American mathematics 1890–1913: catching up to Europe, by Steve Batterson. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):200–201, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1489618.

Abrams:2020:IPL

[Abr20] Ellen Abrams. ‘An inalienable prerogative of a liberated spirit’: postulating American mathematics. British Journal for the History of Mathematics, 35(3):225–245, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/26375451.2020.1805934.

Anderson:2013:MWD

[AC13a] Ian Anderson and Tony Crilly. The mathematician who drove whist forward: William Henry Whitfeld (1856–1915). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28 (3):132–142, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.788891.

Anderson:2013:RRA

[AC13b] Ian Anderson and Tony Crilly. Robert Richard Anstice (1813– 1853): a Hertfordshire bicentenary. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):75–83, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.727146.

Acheson:2023:TSD

[Ach23]  David Acheson.  Thomas

Simpson and Dido’s problem. British Journal for the History of Mathematics, 38(3):189–193, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2250980.

Almira:2019:WDH

[ACO19] José María Almira, José Ángel Cid, and Julio Ostalé. When did Hermann Weyl pass away? British Journal for the History of Mathematics, 34(1):60–63, 2019. CODEN ???? ISSN 2637-5494. URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2018.1555926.

Aydin:2016:AHM

[ADD16] Emin Aydın, Ali Delice, and Derya Demiroğlu. An analysis of history of mathematics research literature in Turkey: the mathematics education perspective. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):215–229, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1190201.

Arzarello:2009:NN

[AFGM09] Ferdinando Arzarello, Fulvia Furinghetti, Livia Giacardi, and Marta Menghini. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):48–49, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802566069.

Aggarwal:2007:MBI

[Agg07] Abhilasha Aggarwal. Mathematical books for and in India in the nineteenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):11–21, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430601148671.

Ackerberg-Hastings:2008:JPB

[AH08] Amy Ackerberg-Hastings. John Playfair on British decline in mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):81–95, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802037830.

Alassi:2020:JBA

[Ala20] Sepideh Alassi. Jacob Bernoulli’s analyses of the Funicularia problem. British Journal for the History of Mathematics, 35(2):137–161, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1741302.

Aldrich:2021:MWB

[Ald21] John Aldrich. Mathematical women in the British Isles 1878–1940: using the Davis archive. British Journal for the History of Mathematics, 36(3):210–218, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2021.2003657.

Amson:2008:GML

[Ams08] John Ceres Amson. Gregory’s meridian line of 1673– 74: a St. Andrews detective story. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):58–72, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802019804.

Anderson:2009:CMH

[And09] David Anderson. The contribution of M H A Newman and his mathematicians to the creation of the Manchester ‘Baby’. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):27–39, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802554503.

Anonymous:2004:VS

[Ano04] Anonymous. Volume 19, supplement. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 19(S1):1–19, 2004.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/09629410408000021.

Anonymous:2005:VNBa

[Ano05a]  Anonymous.  Volume 20,

number 1, 2005 (Bulletin 4). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 20(1):1–58, 2005. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/09629410508000011.

Anonymous:2005:VNBb

[Ano05b]  Anonymous.  Volume 20,

number 2, 2005 (Bulletin 5). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 20(2):1–66, 2005. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/09629410508000021.

Anonymous:2005:VNBc

[Ano05c]  Anonymous.  Volume 20,

number 3, 2005 (Bulletin 6). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 20(3):1–46, 2005. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/09629410508000031.

Anonymous:2006:BMa

[Ano06a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):51–58, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600566675.

Anonymous:2006:BMb

[Ano06b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):132–142, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804514.

Anonymous:2006:BMc

[Ano06c] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):230–233, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600964524.

Anonymous:2006:E

[Ano06d] Anonymous. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):1, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600566477.

Anonymous:2006:NNa

[Ano06e] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):152–154, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600804993.

Anonymous:2006:NNb

[Ano06f] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):236–238, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600964557.

Anonymous:2006:OM

[Ano06g]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):234–235, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600964532.

Anonymous:2006:OMEa

[Ano06h] Anonymous. Other meetings and events. BSHM Bulletin:

Journal of the British Society for the History of Mathematics, 21(1):59–72, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600566758.

Anonymous:2006:OMEb

[Ano06i] Anonymous. Other meetings and events. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):143–151, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804951.

Anonymous:2006:RPa

[Ano06j]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):67–72, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600888160.

Anonymous:2006:RPb

[Ano06k]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):158–161, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/17498430600805065.

Anonymous:2006:RPc

[Ano06l]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):239–244, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600964565.

Anonymous:2006:R

[Ano06m] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21 (2):155–157, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600805032.

Anonymous:2007:BMa

[Ano07a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):48–50, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430601149232.

Anonymous:2007:BMb

[Ano07b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):133–138, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701340863.

Anonymous:2007:BMc

[Ano07c] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):185–191, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701581060.

Anonymous:2007:LRC

[Ano07d] Anonymous. A letter in response to Chris Pritchard’s article on Galton’s quincunx. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):47, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430601149224.

See [Pri06, Pri08].

Anonymous:2007:NNa

[Ano07e] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):60–65, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430601149281.

Anonymous:2007:NNb

[Ano07f] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):140–141, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701340897.

Anonymous:2007:NNc

[Ano07g] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):194–195, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701581136.

Anonymous:2007:OMa

[Ano07h]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):51–59, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430601149257.

Anonymous:2007:OMb

[Ano07i]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):139, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339733.

Anonymous:2007:OMc

[Ano07j]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):192–193, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701581078.

Anonymous:2007:PMH

[Ano07k]  Anonymous.  Perceptions

of mathematics and its history. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):77, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701356281.

Anonymous:2007:RPa

[Ano07l]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):72–76, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/17498430601149422.

Anonymous:2007:RPb

[Ano07m]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):146–149, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701339758.

Anonymous:2007:RPc

[Ano07n]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):201–205, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701581110.

Anonymous:2007:Ra

[Ano07o] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1): 66–71, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/17498430601149414.

Anonymous:2007:Rb

[Ano07p] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22 (2):142–145, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339741.

Anonymous:2007:Rc

[Ano07q] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22 (3):196–200, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701581086.

Anonymous:2008:BMa

[Ano08a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):42–47, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701790455.

Anonymous:2008:BMb

[Ano08b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):109–118, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802025892.

Anonymous:2008:BMc

[Ano08c] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):181–188, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304057.

Anonymous:2008:NNa

[Ano08d] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):50, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701790489.

Anonymous:2008:NNb

[Ano08e] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):119–120, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802037855.

Anonymous:2008:NNc

[Ano08f] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):189–194, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304073.

Anonymous:2008:OM

[Ano08g]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):48–49, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701790463.

Anonymous:2008:RPa

[Ano08h]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):51–55, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701799191.

Anonymous:2008:RPb

[Ano08i]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):123–127, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802025934.

Anonymous:2008:RPc

[Ano08j]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):200–204, 2008.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304115.

Anonymous:2008:Ra

[Ano08k] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23 (2):121–122, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802025926.

Anonymous:2008:Rb

[Ano08l] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23 (3):195–199, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304099.

Anonymous:2009:BMa

[Ano09a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):40–47, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802551939.

Anonymous:2009:BMb

[Ano09b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):185–188, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903088518.

Anonymous:2009:NNa

[Ano09c] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):119–120, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430902848680.

Anonymous:2009:NNb

[Ano09d] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):189–190, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903088534.

Anonymous:2009:RPa

[Ano09e]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):62–66, 2009.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802551988.

Anonymous:2009:RPb

[Ano09f]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):131–134, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/17498430902820911.

Anonymous:2009:RPc

[Ano09g]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):200–202, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/17498430903049544.

Anonymous:2010:BMa

[Ano10a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):119–121, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498431003659869.

Anonymous:2010:BMb

[Ano10b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):179–180, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.497293.

Anonymous:2010:NNa

[Ano10c] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):42–43, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903384081.

Anonymous:2010:NNb

[Ano10d] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):129–130, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498431003659877.

Anonymous:2010:NNc

[Ano10e] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):190, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2010.503689.

Anonymous:2010:OMa

[Ano10f]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):41, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903384073.

Anonymous:2010:OMb

[Ano10g]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):122–126, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498431003664463.

Anonymous:2010:RP

[Ano10h]  Anonymous.  Recent pub-

lications. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):52–55, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/abs/
                                        10.1080/17498430903407932.

Anonymous:2011:BMa

[Ano11a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):53–57, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.529042.

Anonymous:2011:BMb

[Ano11b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):178–182, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.587151.

Anonymous:2012:BMa

[Ano12a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):56–59, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.625343.

Anonymous:2012:BMb

[Ano12b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(3):181–183, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.687497.

Anonymous:2012:NN

[Ano12c] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):70–71, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.628514.

Anonymous:2012:OM

[Ano12d]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):60–61, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.625345.

Anonymous:2013:BMa

[Ano13a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1):52–53, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.729921.

Anonymous:2013:BMb

[Ano13b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):156–163, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.806427.

Anonymous:2013:NN

[Ano13c] Anonymous. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1):59–63, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.729922.

Anonymous:2013:OM

[Ano13d]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):164–165, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.806430.

Anonymous:2013:R

[Ano13e] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28 (3):166–169, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.809504.

Anonymous:2014:BMa

[Ano14a] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):68–70, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.853236.

Anonymous:2014:BMb

[Ano14b] Anonymous. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):134–136, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.894240.

Anonymous:2014:EB

[Ano14c] Anonymous. Editorial board. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):ebi, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.964499.

Anonymous:2014:OM

[Ano14d]  Anonymous.  Other meet-

ings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):137–138, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.894667.

Anonymous:2014:Ra

[Ano14e] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1): 71–79, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430.2014.
855971.

Anonymous:2014:Rb

[Ano14f] Anonymous. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29 (2):139–145, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.894323.

Anonymous:2015:EB

[Ano15] Anonymous. Editorial board.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):ebi, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1102417.

Anonymous:2017:E

[Ano17a] Anonymous. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):185, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1356042.

Anonymous:2017:MET

[Ano17b]  Anonymous.  Mathematics

emerging: A tribute to Jackie Stedall and her influence on the history of mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):100–111, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1220051.

Anonymous:2017:NSM

[Ano17c] Anonymous. News sheet march 2017. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):183, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1314099.

Anonymous:2018:BMP

[Ano18a] Anonymous. BSHM meeting: The practice of mathematics from the Renaissance to the Nineteenth Century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):68–69, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2017.1384203.

Anonymous:2018:RP

[Ano18b]  Anonymous.  Research in

progress. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):144–146, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1450035.

Anonymous:2019:TFE

[Ano19a] Anonymous. 2019 Taylor and Francis Early Career Research Prize. British Journal for the History of Mathematics, 34(1): 69, 2019. CODEN ???? ISSN 2637-5494. URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2018.1542198.

Anonymous:2019:BMN

[Ano19b] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 34(3):200–210, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1654298.

Anonymous:2019:BMR

[Ano19c] Anonymous. BSHM Meeting reports. British Journal for the History of Mathematics, 34(2):125–132, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1591703.

Anonymous:2019:HCC

[Ano19d]  Anonymous.  The history

of cryptography and codes. British Journal for the History of Mathematics, 34(1):71–72, 2019. CODEN ???? ISSN 2637-5494. URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1542200.

Anonymous:2019:HWB

[Ano19e] Anonymous. History workshop at the British Mathematical Colloquium. British Journal for the History of Mathematics, 34(1):73–76, 2019. CODEN ???? ISSN 2637-

                                        5494.  URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1556462.

Anonymous:2019:MP

[Ano19f] Anonymous. Mathematics and patronage. British Journal for the History of Mathematics, 34(1):77–78, 2019. CO-DEN ???? ISSN 2637-

                                        5494.  URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1556460.

Anonymous:2019:P

[Ano19g] Anonymous. Prizes. British Journal for the History of Mathematics, 34(1):70, 2019. CODEN ???? ISSN 2637-

                                        5494.  URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1545893.

Anonymous:2020:BMNa

[Ano20a] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 35(2):180–186, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1745494.

Anonymous:2020:BMNb

[Ano20b] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 35(3):256–262, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1811579.

Anonymous:2020:P

[Ano20c] Anonymous. Prizes. British Journal for the History of Mathematics, 35(1):105–106, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/26375451.2020.
1702678.

Anonymous:2020:WTF

[Ano20d] Anonymous. Winner of the 2018–19 Taylor and Francis Early Career Prize. British Journal for the History of Mathematics, 35(2):187, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1745495.

Anonymous:2021:BMNa

[Ano21a] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 36(1):59–63, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1853258.

Anonymous:2021:BMNb

[Ano21b] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 36(2):140–149, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1911437.

Anonymous:2021:P

[Ano21c] Anonymous. Prizes. British Journal for the History of Mathematics, 36(1):64, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1851443.

Anonymous:2021:SN

[Ano21d] Anonymous. Society news. British Journal for the History of Mathematics, 36(3): 222, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/

full/10.1080/26375451.2021.
1989756.

Anonymous:2022:BMNa

[Ano22a] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 37(1):86–95, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2039514.

Anonymous:2022:BMNb

[Ano22b] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 37(2):164–169, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2086386.

Anonymous:2022:BMNc

[Ano22c] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 37(3):258–265, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2135064.

Anonymous:2022:P

[Ano22d] Anonymous. Prizes. British Journal for the History of Mathematics, 37(1):99, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/26375451.2022.2039519.

Anonymous:2023:BMG

[Ano23a] Anonymous. Between music and geometry: a proposal for the early intended application of Euclid’s Elements book X. British Journal for the History of Mathematics, 38(2):69–96, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2197351.

Anonymous:2023:BMR

[Ano23b] Anonymous. British mathematical reformers in the nineteenth century: motivations and methods: Winner of the BSHM undergraduate essay prize 2022. British Journal for the History of Mathematics, 38(2):159–167, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2222414.

Anonymous:2023:BMNa

[Ano23c] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 38(1):58–66, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2180284.

Anonymous:2023:BMNb

[Ano23d] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 38(2):158, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2224600.

Anonymous:2023:CIA

[Ano23e] Anonymous. The construction of the integral for the arc length of a curve based on van Heuraet and Fermat’s works. British Journal for the History of Mathematics, 38(1):41–54, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2168880.

Anonymous:2023:CRD

[Ano23f] Anonymous. A critical rendition to the development of mathematics education in Nepal: an anticolonial proposal. British Journal for the History of Mathematics, 38(1):3–23, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2109832.

Anonymous:2023:E

[Ano23g] Anonymous. Editorial. British Journal for the History of Mathematics, 38(1):1–2, 2023.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2184033.

Anonymous:2023:GML

[Ano23h] Anonymous. Graphs in the 1680s: Martin Lister, Robert Plot, William Molyneux and John Warner. British Journal for the History of Mathematics, 38(2):97–106, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2224132.

See addendum [Ano24a].

Anonymous:2023:IAP

[Ano23i]  Anonymous.  Infinite an-

alytical procedures for the computation of logarithms in works by Benito Bails (1731– 1797). British Journal for the History of Mathematics, 38(2):107–140, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2186648.

Anonymous:2023:DSD

[Ano23j] Anonymous. Obituary David Singmaster (December 1938 to 13 February 2023). British Journal for the History of Mathematics, 38(2):168–173, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/26375451.2023.
2225798.

Anonymous:2023:PA

[Ano23k]  Anonymous.  Prizes and

awards. British Journal for the History of Mathematics, 38(1):67–68, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2193113.

Anonymous:2023:RME

[Ano23l] Anonymous. Reading mathematics in Early Modern Europe: Studies in the production, collection, and use of mathematical books. British Journal for the History of Mathematics, 38(1):55–57, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/26375451.2023.
2169567.

Anonymous:2023:SUE

[Ano23m] Anonymous. Schools and Undergraduate Essay Prize Winners 2023. British Journal for the History of Mathematics, 38(3):233, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2275111.

Anonymous:2023:SAK

[Ano23n] Anonymous. The sine anecdote in Kovalevskaya’s memoirs. British Journal for the History of Mathematics, 38(1):24–40, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2135298.

Anonymous:2023:TLG

[Ano23o] Anonymous. Through the looking glass, and what algebra found there: historically informed conceptual metaphors of algebraic substitution and Gaussian elimination. British Journal for the History of Mathematics, 38(2):141–157, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2215650.

Anonymous:2024:AGM

[Ano24a]  Anonymous.  Addendum

to ‘Graphs in the 1680s: Martin Lister, Robert Plot, William Molyneux and John Warner’. British Journal for the History of Mathematics, 39(1):70, 2024. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2317697.

See [Ano23h].

Anonymous:2024:BSH

[Ano24b]  Anonymous.  British Soci-

ety for the History of Mathematics prizes and awards. British Journal for the History of Mathematics, 39(1): 74–75, 2024. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/26375451.2024.
2319966.

Anonymous:2024:BMN

[Ano24c] Anonymous. BSHM meeting news. British Journal for the History of Mathematics, 39(1):71–72, 2024. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2317694.

Anonymous:2024:BNP

[Ano24d] Anonymous. BSHM Neumann Prize Winner 2023. British Journal for the History of Mathematics, 39(1):73, 2024. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2319468.

Astarte:2020:DCM

[Ast20]  Troy Astarte.  The dis-

crete charm of the machine: why the world became digital. British Journal for the History of Mathematics, 35(1):89–91, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1681827.

Atzema:2015:BEF

[Atz15] Eisso J. Atzema. From Brahmagupta to Euler: on the formula for the area of a cyclic quadrilateral. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):20–34, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.942818.

Bailie:2021:QUR

[Bai21] Natasha Bailie. Quantifying the unquantifiable: the role of the mathematicisation of philosophy during the Scottish Enlightenment. British Journal for the History of Mathematics, 36(2):132–139, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2021.1880137.

Ball:2015:TRF

[Bal15]  Derek Ball.  ‘Thick-rinded

fruit of the tree of knowledge’: mathematics education in George Eliot’s novels. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):217–226, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1044161.

Barnett:2009:MGB

[Bar09] Janet Heine Barnett. Mathematics goes ballistic: Benjamin Robins, Leonhard Euler, and the mathematical education of military engineers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):92–104, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430902820887.

Basyal:2020:MPB

[Bas20] Deepak Basyal. A mathematical poetry book from Nepal. British Journal for the History of Mathematics, 35(3):189–206, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1777781.

Bayley:2007:HTS

[Bay07] Mel Bayley. Hard times and statistics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):92–103, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339600.

Bair:2021:PLI

[BBE^{+}21] Jacques Bair, Piotr Błaszczyk, Robert Ely, Mikhail G. Katz, and Karl Kuhlemann. Procedures of Leibnizian infinitesimal calculus: an account in three modern frameworks. British Journal for the History of Mathematics, 36(3):170–209, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1851120.

Bair:2020:CBC

[BBG^{+}20] Jacques Bair, Piotr Błaszczyk, Elías Fuentes Guillén, Peter Heinig, Vladimir Kanovei, and Mikhail G. Katz. Continuity between Cauchy and Bolzano: issues of antecedents and priority. British Journal for the History of Mathematics, 35(3):207–224, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1770015.

Beery:2009:FFN

[Bee09] Janet L. Beery. Formulating figurate numbers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2): 78–91, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430902820879.

Beeley:2017:PAJ

[Bee17] Philip Beeley. ‘To the publike advancement’ John Collins and the promotion of mathematical knowledge in Restoration England. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):61–74, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1278829.

Beeley:2020:TGA

[Bee20] Philip Beeley. ‘There are great alterations in the geometry of late’. The rise of Isaac Newton’s early Scottish circle. British Journal for the History of Mathematics, 35(1):3–24, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1701862.

Bellos:2010:MPF

[Bel10] David Bellos. Mathematics, poetry, fiction: the adventure of the Oulipo. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):104–118, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903489237.

Bellhouse:2014:DN

[Bel14] David R. Bellhouse. The deification of Newton in 1711. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):98–110, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.845931.

Bellhouse:2015:MEE

[Bel15a] David R. Bellhouse. Mathematicians and the early English life insurance industry. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):131–142, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1004895.

Bellissima:2015:PVE

[Bel15b] Fabio Bellissima. Propositions VIII.4–5 of Euclid’s Elements and the compounding of ratios on the monochord. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):183–199, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1044697.

Berenguer:2021:IDC

[Ber21] Joaquim Berenguer. Introducing differential calculus in Spain: The fluxion of the product and the quadrature of curves by Tomàs Cerdà. British Journal for the History of Mathematics, 36(1):23–49, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1829864.

Barrow-Green:2006:MNA

[BG06] June Barrow-Green. ‘Much necessary for all sortes of men’: 450 years of Euclid’s Elements in English. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):2–25, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600566527.

Barrow-Green:2010:EE

[BG10] June Barrow-Green. Euler as an educator. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):10–22, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903321141.

Bruno:2011:RPH

[BGI11] Giuseppe Bruno, Andrea Genovese, and Gennaro Improta. Routing problems: a historical perspective. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):118–127, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.560092.

Bruno:2014:HPL

[BGI14] Giuseppe Bruno, Andrea Genovese, and Gennaro Improta. A historical perspective on location problems. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):83–97, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.889266.

Biggs:2009:MCE

[Big09]  Norman Biggs.  Mathe-

matics of currency and exchange: arithmetic at the end of the thirteenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):67–77, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430902820861.

Biggs:2013:THC

[Big13] Norman Biggs. Thomas Harriot on continuous compounding. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):66–74, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.721331.

Biggs:2017:MSC

[Big17] Norman Biggs. More seventeenthcentury networks. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1): 30–39, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430.2016.
1225781.

Biggs:2018:GSG

[Big18] Norman Biggs. Game, set, and graph. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):166–178, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1504458.

Biggs:2019:BRD

[Big19]  Norman Biggs.  Book re-

view: Dan Bouk, How our days became numbered: risk and the rise of the statistical individual. British Journal for the History of Mathematics, 34(2):123–124, 2019.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1568106.

Biggs:2021:DCR

[Big21]  Norman Biggs.  Decoding

chancery records from the 1240s. British Journal for the History of Mathematics, 36(1):1–12, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1834796.

Biggs:2022:WME

[Big22] Norman Biggs. William Morgan, Eighteenth-century actuary, mathematician and radical. British Journal for the History of Mathematics, 37(1):81–82, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.2009720.

Bjarnadottir:2009:PR

[Bja09]  Kristín Bjarnadóttir.  A

puzzle rhyme from 1782. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):12–19, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802551921.

Bingham:2022:LAM

[BK22] N. H. Bingham and W. J.

Krzanowski. Linear algebra and multivariate analysis in statistics: development and interconnections in the twentieth century. British Journal for the History of Mathematics, 37(1):43–63, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2045811.

Blackburn:2008:MWS

[Bla08]  Simon R. Blackburn.  A

mathematical walk in Surrey. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):178–180, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304040.

Boag:2007:LM

[Boa07] Elizabeth Boag. Lattice multiplication. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):182–184, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701581045.

Bowen:2015:BRI

[Bow15] Jonathan P. Bowen. Book review: It began with Babbage: the genesis of computer science, by Subrata Dasgupta. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):263–265, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1036336.

Boyd:2024:NAP

[Boy24] T. James M. Boyd. Note added in proof. British Journal for the History of Mathematics, 39(1):62–63, 2024. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2326795.

Bradley:2013:HBG

[Bra13]  Robert E. Bradley.  De

l’Hôpital, Bernoulli, and the genesis of Analyse des infiniment petits. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1):16–24, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.729916.

Brantley:2015:FCS

[Bra15] Kristina Leifeste Brantley. A forgotten contrivance: a study of the diagonal scale and its appearance in mathematics texts from 1714 to the present. BSHM Bulletin:

Journal of the British Society for the History of Mathematics, 30(1):50–66, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.917391.

Boyd:2024:STL

[BRO24] T. James M. Boyd, Gareth Roberts, and Alwyn R. Owens. Stability in theory, in the laboratory and in the air: William Ellis Williams’ campaign for proof positive (1904– 1914). British Journal for the History of Mathematics, 39(1):36–61, 2024. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2312323.

Bruneau:2020:CMN

[Bru20]  Olivier Bruneau.  Colin

Maclaurin (1698–1746): a Newtonian between theory and practice. British Journal for the History of Mathematics, 35(1):52–62, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1701859.

Burn:2006:NBL

[Bur06] Bob Burn. The Neil Bibby Lecture 2006: From Archimedes to limits: understanding real analysis. BSHM Bulletin:

Journal of the British Society for the History of Mathematics, 21(2):73–85, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600803276.

Burn:2007:GP

[Bur07] Bob Burn. Geometric progressions. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):183–184, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701581052.

Burn:2013:REE

[Bur13] Bob Burn. Root 2: the early evidence and later conjectures. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):114–120, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.793563.

Burn:2015:LE

[Bur15] Bob Burn. Letter to the editor. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):90–91, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2014.993133.

Burn:2016:ETR

[Bur16] Bob Burn. Early tables resembling those of natural logarithms. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):112–122, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1116052.

Bussotti:2018:DFA

[Bus18] Paolo Bussotti. Book review:

Discovery of the First Asteroid, Ceres, by Clifford Cunningham. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):57–59, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1286559.

Bullock:2022:WWL

[BWH22] John D. Bullock, Ronald E.

Warwar, and H. Bradford Hawley. Why was Leonhard Euler blind? British Journal for the History of Mathematics, 37(1):24–42, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2052493.

Caglayan:2016:ELH

[Cag16] Günhan Caglayan. Exploring the lunes of Hippocrates in a dynamic geometry environment. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):144–153, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1122301.

Caglayan:2019:TPN

[Cag19] Günhan Caglayan. Theory of polygonal numbers with Cuisenaire rods manipulatives: Understanding Theon of Smyrna’s arithmetic in a history of mathematics classroom. British Journal for the History of Mathematics, 34(1):12–22, 2019. CODEN ???? ISSN 2637-5494. URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2018.1539375.

Calinger:2018:BRE

[Cal18] Ronald Calinger. Book review: The early period of the calculus of variations, by Paolo Freguglia and Mariano Giaquinta. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):138–139, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1419700.

Cameron:2023:GTA

[Cam23] Peter Cameron. Graph theory in America: the first hundred years. British Journal for the History of Mathematics, 38(3):229–230, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2248446.

Carman:2018:BRP

[Car18] Christián C. Carman. Book review: A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World, by Alexander Jones. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):189–190, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1457342.

Cooper:2013:FLH

[CC13] Stephen Cooper and Ashley Cooper. The father of local history and the duplication of the exchequer, 1480. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):121–123, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.802113.

Clark:2016:PMD

[CH16] Kathleen M. Clark and Emmet P. Harrington. The Paul A. M. Dirac papers at Florida State University: a search for informal mathematical investigations. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):205–214, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1175143.

Christiansen:2009:BMH

[Chr09] Andreas Christiansen. Bernt Michael Holmboe (1795–1850) and his mathematics textbooks. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):105–113, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430902820895.

Campbell-Kelly:2007:WBW

[CK07] Martin Campbell-Kelly. From the world brain to the World Wide Web Annual Gresham College BSHM Lecture. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):1–10, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430601148614.

Clark:2010:CLH

[Cla10] Kathleen M. Clark. Connecting local history, ancient history, and mathematics: the eustis elementary school pilot project. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):132–143, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903449199.

Clark:2020:BRW

[Cla20] Kathleen M. Clark. Book review: Women who count: Honoring African American women mathematicians, Shelly M. Jones, American Mathematical Society, 2020, xiii + 138pp., $15.00, ISBN 978-1-4704-4889-9. British Journal for the History of Mathematics, 35(3):253–255, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1778282.

Cretney:2012:R

[CMBV12] Rosanna Cretney, Indra Kagis McEwen, Melanie Bayley, and Alexei Volkov. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(3):184–191, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.687678.

Craik:2011:SUD

[CO11] Alex D. D. Craik and John J. O’Connor. Some unknown documents associated with William Wallace (1768–1843). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):17–28, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2010.503555.

Coates:2008:EWZ

[Coa08] John Coates. Euler’s work on zeta and L-functions and their special values. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):37–41, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701790422.

Clark:2008:AAM

[CR08] Kathleen Clark and Eleanor Robson. Ancient accounting in the modern mathematics classroom. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):129–142, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430802304008.

Craik:2006:AMC

[Cra06] Alex Craik. Arthur Mee’s Children’s encyclopedia revisited. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):122–126, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804449.

Craik:2010:WWC

[Cra10] Alex D. D. Craik. William Wallace’s chorograph (1839): a rare mathematical instrument. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):23–31, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903321158.

Craik:2012:PLA

[Cra12] Alex D. D. Craik. The popular lectures and addresses of William Thomson, Baron Kelvin of Largs (1824–1907). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):50–55, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.619085.

Craik:2013:STK

[Cra13] Alex D. D. Craik. In search of Thomas Knight: Part

2. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):124–131, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.783970.

Craik:2015:BKJ

[Cra15] Alex D. D. Craik. A book for the king: John Geddy’s Methodi sive compendii mathematici (1586). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):113–130, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1004506.

Crampin:2016:NSR

[Cra16] Stephanie Crampin. A note on some recent mathematical histories/A curious history of mathematics, by Joel Levy/ The universe in zero words, by Dana Mackenzie/17 equations that changed the world, by Ian Stewart/Maths in 100 key breakthroughs, by Richard Elwes. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):83–86, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1080442.

Craik:2017:ESP

[Cra17]  Alex D. D. Craik.  An

early Scottish pamphlet on hydraulics and pneumatics: William Welwood’s De aqua in altum per fistulas plumbeas facile exprimenda apologia demonstrativa (1582). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (2):113–124, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1278657.

Craik:2020:GSN

[Cra20a] Alex D. D. Craik. George Sinclair’s neglected treatises: some influences and reactions. British Journal for the History of Mathematics, 35(1):43–51, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1673101.

Craik:2020:HPH

[Cra20b] Alex D. D. Craik. Henry Parr Hamilton (1794–1880) and analytical geometry at Cambridge. British Journal for the History of Mathematics, 35(2):162–170, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1735824.

Crilly:2015:BLM

[Cri15] Tony Crilly. BSHM–LMS De Morgan Day. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):253–255, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1055089.

Crilly:2016:BCM

[Cri16] Tony Crilly. BSHM Christmas meeting Birmingham, 5 December 2015. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):246–249, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1215858.

Crilly:2018:BRT

[Cri18a] Tony Crilly. Book review: Ten great ideas about chance, by Persi Diaconis and Brian Skyrms. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):197–199, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1478532.

Crilly:2018:WBP

[Cri18b]  Tony Crilly.  What be-

came of Paul Dirac’s classmate? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):179–188, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1518508.

Crilly:2020:BRP

[Cri20a]  Tony Crilly.  Book re-

view: Pythagoras’ legacy, Marcel Danesi, Oxford University Press, 2020, 167 pp., £25, ISBN 978-0-19-885224-

7. British Journal for the History of Mathematics, 35 (3):246–247, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1765504.

Crippa:2020:BUP

[Cri20b] Davide Crippa. Beating untrodden paths: James Gregory and his Italian readers. British Journal for the History of Mathematics, 35(1):25–42, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1701860.

Camunez-Ruiz:2019:JCS

[CRPH19] José Antonio Camúñez-Ruiz and María Dolores Pérez-Hidalgo. Juan Caramuel (1606–1682) and the Spanish version of the Passedix game. British Journal for the History of Mathematics, 34(3):143–154, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1630986.

Cuomo:2012:EAG

[Cuo12] Serafina Cuomo. Exploring ancient Greek and Roman numeracy. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):1–12, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.618101.

Dauben:2020:BRH

[Dau20] Joseph W. Dauben. Book review: Heavenly numbers. Astronomy and authority in early imperial China, Christopher Cullen, Oxford University Press, 2017, 448pp, £74, ISBN 978-0-1987-3311-

9. British Journal for the History of Mathematics, 35 (3):247–251, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1780837.

Davis:2017:MWC

[Dav17] A. E. L. Davis. Mathematical women — creating historical resources. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):254–256, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1318248.

Davis:2018:WMO

[Dav18] A. E. L. Davis. Woman into mathematician: The opening of university mathematical education to women in the British Isles: a prosopographical note. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):29–32, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1308126.

Dawson:2018:BRG

[Daw18] John W. Dawson, Jr. Book review: The Great Formal Machinery Works: Theories of Deduction and Computation at the Origins of the Digital Age, by Jan Von Plato. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):196–197, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1472471.

deBruycker:2009:AHC

[dB09]  Angelo de Bruycker.  ‘To

the adornment and honour of the city’: the mathematics course of the Flemish Jesuits in the seventeenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):135–146, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903045880.

Diniz:2021:DGJ

[DB21] Marcio Alves Diniz and David Richard Bellhouse. David Gregory, John Arbuthnot and their roles in the early development of probability in Great Britain. British Journal for the History of Mathematics, 36(1):13–22, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1847505.

Dean:2011:IPT

[Dea11] Amanda Saxon Dean. An investigation of pedagogical techniques in Descartes’ La géométrie. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):167–177, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.588586.

Decesare:2011:WLP

[Dec11] Richard Decesare. William Ludlam: portrait of an eighteenth-century mathematician. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):105–117, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.529043.

DeCesare:2016:RPA

[DeC16] Richard DeCesare. Robert Patterson: American ‘revolutionary’ mathematician. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):31–46, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1035588.

Dejic:2014:HOS

[Dej14] Mirko Dejić. How the old Slavs (Serbs) wrote numbers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):2–17, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.805559.

Denniss:2006:ATP

[Den06] John Denniss. Arithmetical textbooks 1478 to 1886: a progression? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):26–33, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600566550.

Denniss:2015:BRR

[Den15] John Denniss. Book review: Rewriting the history of school mathematics in North America 1607–1861: the central role of cyphering books, by Nerida Ellerton and M. A. (Ken) Clements. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):82–83, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.919434.

Dewar:2017:WMC

[Dew17] Jacqueline M. Dewar. Women and mathematics: a course and a scholarly investigation. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):246–253, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1319160.

Domingues:2014:RJA

[Dom14] João Caramalho Domingues.

The repercussion of José Anastácio da Cunha in Britain and USA in the nineteenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):32–50, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.802111.

Drucker:2016:LE

[Dru16] Thomas Drucker. Letter to the editor. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):87, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1125708.

Edwards:2006:ECV

[Edw06]  A. W. F. Edwards.  An

eleventh-century Venn diagram. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):119–121, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804407.

Eminger:2012:VUS

[Emi12] Stefanie Eminger. Viribus unitis! shall be our watchword: the first International Congress of Mathematicians, held 9–11 August 1897 in Zurich. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(3):155–168, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.687496.

Enflo:2022:HSS

[EMSS22] P. H. Enflo, M. S. Moslehian, and J. B Seoane-Sepúlveda. A history of solving some famous problems in mathematical analysis. British Journal for the History of Mathematics, 37(1):64–80, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2037358.

Fadlelmula:2015:PST

[Fad15] Fatma Kayan Fadlelmula. Preservice teachers’ point of views about learning history of mathematics: a case study in Turkey. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):243–252, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1035585.

Fairbairn:2017:LJM

[Fai17] Ben Fairbairn. Louis Joel Mordell’s time in London. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (2):160–169, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1233692.

Falconer:2018:BM

[Fal18] Isobel Falconer. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):67–68, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1325298.

Falconer:2020:E

[Fal20a]  Isobel Falconer.  Edito-

rial. British Journal for the History of Mathematics, 35(1):1–2, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1702651.

Falconer:2020:PAA

[Fal20b]  Isobel Falconer.  Professor

Alexander (Alex) D. D. Craik: 1938–2019. British Journal for the History of Mathematics, 35(2):178–179, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1745493.

Falconer:2022:E

[Fal22]  Isobel Falconer.  Edito-

rial. British Journal for the History of Mathematics, 37(2):101–102, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2092827.

Falconer:2024:E

[Fal24]  Isobel Falconer.  Edito-

rial. British Journal for the History of Mathematics, 39(1):1–2, 2024. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2024.2319452.

Falconer:2020:BMN

[FC20] Isobel Falconer and (BSHM Meeting Coordinator:). BSHM meeting news. British Journal for the History of Mathematics, 35(1):94–104, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1702670.

Feather:2018:BRC

[Fea18] Felix Feather. Book review: The Calculus Story: A Mathematical Adventure, by David Acheson. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):192–194, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1472472.

Findlay:2019:BRD

[Fin19] Matthew Findlay. Book review: David Aubin, L’Élite sous la mitraille: les normaliens, les mathématiques et la Grande Guerre, 1900–1925, Éditions Rue d’Ulm, 2018, 374 pp., 25 EUR, ISBN 978-2-7288-0603-4. British Journal for the History of Mathematics, 34(3):197–199, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1617583.

Flood:2006:MVI

[Flo06]  Raymond Flood.  Mathe-

matics in Victorian Ireland. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21 (3):200–211, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600964433.

Flood:2013:JCM

[Flo13] Raymond Flood. James Clerk Maxwell (BSHM–Gresham Lecture 2012). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):84–96, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.755073.

Flood:2016:BGC

[Flo16]  Raymond Flood.  BSHM/

Gresham College meeting: Women in mathematics: A celebration of the bicentenary of Ada Lovelace Gresham College, London, 29 October 2015. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):245–246, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1215857.

Fischer:2014:NRN

[FR14] Joachim Fischer and Bärbel Ruess. Napier revisited, or a new look at the computation of his logarithms. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29 (3):167–183, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.919746.

Frejd:2013:OAT

[Fre13] Peter Frejd. Old algebra textbooks: a resource for modern teaching. BSHM Bulletin:

Journal of the British Society for the History of Mathematics, 28(1):25–36, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.676963.

Friedman:2016:TBG

[Fri16] Michael Friedman. Two beginnings of geometry and folding: Hermann Wiener and Sundara Row. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):52–68, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1045700.

Friedman:2019:MFD

[Fri19] Michael Friedman. Mathematical formalization and diagrammatic reasoning: the case study of the braid group between 1925 and 1950. British Journal for the History of Mathematics, 34(1):43–59, 2019. CODEN ???? ISSN 2637-5494. URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2018.1533298.

Friedman:2022:MRM

[Fri22] Michael Friedman. Metaphorical reactions in 1932: from the mathematical ‘crisis of intuition’ to ‘reconstruction in the exact sciences’. British Journal for the History of Mathematics, 37(2):138–161, 2022.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2052631.

Froggatt:2015:CCJ

[Fro15]  Terry Froggatt.  Counting

and calculation, a journey through practical mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):75–77, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.942819.

Gage:2017:UAN

[Gag17] Joseph Gage. Undergraduate algebra in nineteenth-century Oxford. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):149–159, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1244630.

Gaull:2010:TSB

[Gau10] Marilyn Gaull. From Tristram Shandy to Bertrand Russell: fiction and mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):81–91, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498431003659851.

Grabiner:2011:R

[GDR^{+}11] Judith V. Grabiner, Thomas L.

Drucker, David E. Rowe, Jean-Guy Prévost, Laurent Mazliak, and Jeffrey A. Oaks. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):58–70, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.530165.

Grattan-Guinness:2014:AWS

[GG14a] I. Grattan-Guinness. Addendum on Wussing and Scriba [MR3169973]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):222, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.945067.

Grattan-Guinness:2014:HWC

[GG14b] Ivor Grattan-Guinness. Hans Wussing and Christoph Scriba: an appreciation. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1): 60–67, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430.2013.
845810.

Gow:2006:LWE

[Gow06a] Rod Gow. Letters of William Emerson and Francis Holliday to the publisher, John Nourse. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):40–50, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600566626.

Gow:2006:LWG

[Gow06b] Rod Gow. Life and work of George Salmon (1819– 1904). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):212–218, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600964474.

Gray:2006:OTC

[Gra06] Jeremy Gray. Overstating their case? Reflections on British mathematics in the nineteenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):178–185, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600964391.

Grabiner:2007:WSH

[Gra07] Judith V. Grabiner. Why should historical truth matter to mathematicians? Dispelling myths while promoting maths. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):78–91, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339584.

Gray:2008:SLE

[Gra08] Jeremy Gray. A short life of Euler. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):1–12, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701790398.

Gray:2009:DSC

[Gra09] Shirley B. Gray. A detective story continued. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24 (3):180–184, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903018598.

Gray:2020:DGS

[Gra20] Jeremy Gray. Descriptive geometry, the spread of a polytechnic art: the legacy of Gaspard Monge. British Journal for the History of Mathematics, 35(1):85–87, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1678821.

Goulding:2017:MTH

[GSttmoJS17] Robert Goulding, Matthias Schemmel, and Dedicated to the memory of Jacqueline Stedall. The manuscripts of Thomas Harriot (1560– 1621). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):17–19, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1260792.

Guicciardini:2018:APH

[Gui18] Niccolò Guicciardini. Un altro presente: on the historical interpretation of mathematical texts. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):148–165, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1518844.

Gulizia:2022:KSE

[Gul22] Stefano Gulizia. Kepler’s snow: the epistemic playfulness of geometry in seventeenth-century Europe. British Journal for the History of Mathematics, 37(2):117–137, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2092370.

Haffner:2019:MLI

[Haf19] Emmylou Haffner. From modules to lattices: Insight into the genesis of Dedekind’s Dualgruppen. British Journal for the History of Mathematics, 34(1):23–42, 2019. CO-DEN ???? ISSN 2637-

5494.  URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1555928.

Hartveit:2009:HFG

[Har09] Marit Hartveit. How Flora got her cap: the higher education of women in Edinburgh. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):147–158, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903008391.

Harper:2016:DCR

[Har16] J. F. Harper. Defining continuity of real functions of real variables. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):189–204, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1116053.

Hart:2021:E

[Har21] Sarah Hart. Editorial. British Journal for the History of Mathematics, 36(2):65–66, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/26375451.2021.
                                        1906536.

Heefer:2019:VCM

[Hee19] Albrecht Heefer. Visual culture and mathematics in the early modern period. British Journal for the History of Mathematics, 34(3):194–195, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1653066.

Hollings:2012:MAF

[HG12] Christopher Hollings and Judith V. Grabiner. Michèle Audin, Fatou, Julia, Montel: the Great Prize of mathematical sciences of 1918, and beyond. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(2):127–130, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.658539.

Higgitt:2019:BRE

[Hig19] Rebekah Higgitt. Book review: Elizabethanne Boran and Mordechai Feingold (eds), Reading Newton in early modern Europe. British Journal for the History of Mathematics, 34(1):64–66, 2019. CODEN ???? ISSN 2637-

5494.  URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1542196.

Hildyard:2014:JPM

[Hil14] Daisy Hildyard. John Pell’s mathematical papers and the Royal Society’s English Atlas, 1678–82. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):18–31, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.813735.

Hindle:2024:DTF

[Hin24] Kate Hindle. D’Arcy Thompson on flight. British Journal for the History of Mathematics, 39(1):26–35, 2024. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2312324.

Hitchcock:2021:BRN

[Hit21] Gavin Hitchcock. Book review: New light on George Boole, Desmond MacHale and Yvonne Cohen, Cork University Press, 2018, 492pp, £13.75, ISBN 978-1-78205-290-6. British Journal for the History of Mathematics, 36(1):51–53, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1819740.

Hollings:2017:EME

[HMR17] Christopher Hollings, Ursula Martin, and Adrian Rice. The early mathematical education of Ada Lovelace. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (3):221–234, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1325297.

Hollings:2006:HSI

[Hol06] Christopher David Hollings. The history of the 2-, 4- and 8-square identities. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21 (2):111–118, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804381.

Hollings:2009:AKS

[Hol09] Christopher Hollings. Anton Kazimirovich Suschkewitsch (1889–1961). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):172–179, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903008417.

Hollings:2014:ICR

[Hol14] Christopher Hollings. Investigating a claim for Russian priority in the abstract definition of a ring. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):111–119, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.858202.

Hollings:2015:BRM

[Hol15] Christopher Hollings. Book review: The Moscow Pythagoreans: Mathematics, Mysticism, and Anti-Semitism in Russian Symbolism, by Ilona Svetlikova. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):85–86, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.954356.

Hollings:2016:TMM

[Hol16a]  Christopher Hollings.  A

tale of mathematical mythmaking: E. T. Bell and the ‘arithmetization of algebra’. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):69–80, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1048640.

Hollings:2016:HNT

[Hol16b] Christopher D. Hollings. The history of number theory, Birkbeck, 21 May 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31 (3):250–252, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1215856.

Hollings:2017:MBM

[Hol17] Christopher D. Hollings. ‘Mathematical biography: A MacTutor celebration’. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):265, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1244749.

Hollings:2018:BRI

[Hol18a] Christopher Hollings. Book review: Images of Italian Mathematics in France: The Latin sisters, from Risorgimento to fascism, by Frédéric Brechenmacher et al. (eds). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33 (2):139–141, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1394714.

Hollings:2018:CAN

[Hol18b] Christopher D. Hollings. Book review: The case of Academician Nikolai Nikolaevich Luzin, by Sergei S. Demidov and Boris V. Lëvshin, translated from the Russian by Roger Cooke. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):59–61, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1244748.

Hollings:2020:RPM

[Hol20] Christopher D. Hollings. A richer picture of mathematics: the Göttingen tradition and beyond. British Journal for the History of Mathematics, 35(1):87–89, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1664134.

Hollings:2021:PMN

[Hol21] Christopher D. Hollings. Peter M Neumann OBE (1940– 2020). British Journal for the History of Mathematics, 36(2):67–75, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1919472.

Hunt:2015:DWS

[HS15] James L. Hunt and John Sharp. Decoding William Scrots’ anamorphic portrait of Edward VI. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):200–216, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1041072.

Huylebrouck:2015:OAL

[Huy15]  Dirk Huylebrouck.  Ob-

servations about Leonardo’s drawings for Luca Pacioli. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30 (2):102–112, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.945066.

Jagger:2006:WHB

[Jag06] Graham Jagger. The will of Henry Briggs. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):127–131, 2006.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804464.

James:2011:RNE

[Jam11] Kathryn James. Reading numbers in early modern England. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):1–16, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.504529.

Jankvist:2008:TMH

[Jan08] Uffe Thomas Jankvist. A teaching module on the history of public-key cryptography and RSA. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):157–168, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802304032.

Jankvist:2014:HTM

[Jan14] Uffe Thomas Jankvist. A historical teaching module on ‘the unreasonable effectiveness of mathematics’: Boolean algebra and Shannon circuits. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):120–133, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.874869.

Jenkins:2010:MMH

[Jen10] Alice Jenkins. Mathematics and mental health in early nineteenth-century England. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):92–103, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498431003696242.

Jesseph:2015:BRT

[Jes15] Douglas Jesseph. Book review: The tangled origins of the Leibnizian calculus: a case study of a mathematical revolution, by Richard C Brown. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):261–263, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1036335.

Karp:2007:WAM

[Kar07] Alexander Karp. ‘We all meandered through our schooling…’: notes on Russian mathematics education in the early nineteenth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22 (2):104–119, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339634.

Karp:2018:MVS

[Kar18] Alexander Karp. Mark Vygodsky: several episodes from the life of a scholar. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):33–49, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1286558.

Kastanis:2007:APG

[Kas07]  Nikos Kastanis.  Ameri-

can pestalozzianism in Greek mathematical education 1830– 1836. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(2):120–132, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701339642.

Kaye:2006:UHM

[Kay06] David Kaye. Using the history of mathematics in training adult numeracy teachers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):86–89, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600803466.

Kent:2012:ABG

[Ken12] Deborah Kent. Alice Bache Gould: mathematician in search of war work, 1918. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):38–49, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.628477.

Kent:2022:EEE

[Ken22] Deborah Kent. Einstein, Eddington, e o/and the eclipse: Impressões de viagem/travel impressions. British Journal for the History of Mathematics, 37(2):162–163, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2056968.

Kidwell:2022:MIT

[Kid22] Peggy Kidwell. Mathematical instruments from times of crisis. British Journal for the History of Mathematics, 37(2):103–116, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2040897.

Knudsen:2011:BRG

[KJB11] Toke Knudsen, Mikkel Willum Johansen, and James Byrne. Book review: Glen van Brummelen, The mathematics of the heavens and the earth: the early history of trigonometry. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):183–188, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.592478.

Klein:2007:QCU

[Kle07] David Klein. A quarter century of US ‘math wars’ and political partisanship. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):22–33, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430601148762.

Kruger:2010:LES

[Krü10] Jenneke Krüger. Lessons from the early seventeenth century for mathematics curriculum design. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):144–161, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903584136.

Kosmann-Schwarzbach:2015:WMF

[KS15] Yvette Kosmann-Schwarzbach. Women mathematicians in France in the mid-twentieth century. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):227–242, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.976804.

Kaufholz-Soldat:2017:FHM

[KS17] Eva Kaufholz-Soldat. ‘[…] the first handsome mathematical lady I’ve ever seen!’ On the role of beauty in portrayals of Sofia Kovalevskaya. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (3):198–213, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1318249.

Kourkoulos:2010:HSU

[KT10] Michael Kourkoulos and Constantinos Tzanakis. History, and students’ understanding of variance in statistics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):168–178, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430903526384.

Kunkel:2007:TPA

[Kun07] Paul Kunkel. The tangency problem of Apollonius: three looks. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(1):34–46, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430601148911.

Lawrence:2006:MGY

[Law06] Snezana Lawrence. Maths is good for you: web-based history of mathematics resources for young mathematicians (and their teachers). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):90–96, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600803375.

Lawrence:2011:DHB

[Law11] Snezana Lawrence. Dee and his books: lessons from the history of mathematics for primary and middle school teachers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):160–166, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.587150.

Lawrence:2016:BRB

[Law16] Snezana Lawrence. Book review: Birth of a theorem: a mathematical adventure, by Cédric Villani (trans Malcolm DeBevoise, ill Claude Gongard). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):263–265, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1208394.

Lawrence:2018:WTS

[Law18] Snezana Lawrence. “What the tortoise said to Achilles”: Lewis Carroll’s paradox of inference, special issue of: The Carrollian: The Lewis Carroll Journal, edited by Amirouche Moktefi and Francine F Abeles. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33 (3):194–195, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1472470.

Leddy:2016:MCD

[Led16a] Dorothy Leddy. Mathematical communication during the Cold War: Mathematical Institute, Oxford, 8 July 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):257–260, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1220065.

Leddy:2016:MER

[Led16b] Dorothy Leddy. Mathematics in the Enlightenment Rewley House, Oxford, 25 June 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):254–256, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1216085.

Leddy:2019:CFW

[Led19] Dorothy Leddy. A common family weakness for statistics: essays on Francis Galton, George Darwin and the normal curve of evolutionary biology. British Journal for the History of Mathematics, 34(3):195–197, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1617585.

Leddy:2021:SN

[Led21]  Dorothy Leddy.  Society

news. British Journal for the History of Mathematics, 36(2):150, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1914424.

Lee:2013:GSA

[Lee13] Alexander Lee. Goals and scope of the Archimedes Palimpsest transcriptions. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1): 1–15, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/17498430.2012.
                                        687974.

Lee:2016:GEM

[Lee16] Peter M. Lee. George Eliot and mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):154, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1162406.

Lewis:2012:PGT

[Lew12] Elizabeth F. Lewis (Née Rudge). P G Tait’s schoolboy introduction to complex numbers. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(3):141–154, 2012. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.687495.

Lewis:2016:PGT

[Lew16]  Elizabeth F. Lewis.  P

G Tait’s statistical models. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31 (2):128–143, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1090096.

Lloyd:2012:HOP

[Llo12] D. R. Lloyd. How old are the Platonic solids? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27 (3):131–140, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.670845.

Lorenat:2012:SSN

[Lor12] Jemma Lorenat. Not set in stone: nineteenth-century geometrical constructions and the Malfatti problem. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27 (3):169–180, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.676962.

MacBeath:2020:AMC

[Mac20] Martin A. MacBeath. Africa and mathematics: from colonial findings back to the Ishango Rods. British Journal for the History of Mathematics, 35(1):91–93, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1684040.

MacanBhaird:2021:BRT

[Mac21] Ciarán Mac an Bhaird. Book review: Thomas Harriot: a life in science, Robyn Arianrhod, Oxford University Press, 2019, 376 pp, £19.99, ISBN 978-0-19027-185-5. British Journal for the History of Mathematics, 36(1):50–51, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1819739.

Magnello:2006:VVM

[Mag06] M. Eileen Magnello. Victorian vital and mathematical statistics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):219–229, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600964508.

Magnello:2012:VSG

[Mag12] M. Eileen Magnello. Victorian statistical graphics and the iconography of Florence Nightingale’s polar area graph. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):13–37, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.618102.

Mallion:2008:CEW

[Mal08] Roger Mallion. A contemporary Eulerian walk over the bridges of Kaliningrad. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):24–36, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701799183.

Mann:2010:SPB

[Man10] Tony Mann. From Sylvia Plath’s The bell jar to the Bad Sex Award: a partial account of the uses of mathematics in fiction. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):58–66, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498431003664455.

Mann:2012:PBC

[Man12] Tony Mann. From prehistoric balls to contemporary fiction: what the history of mathematics has done for me. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(2):73–81, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.660365.

Mann:2013:E

[Man13] Tony Mann. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):113, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.814018.

Mann:2014:Ea

[Man14a] Tony Mann. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):1, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.859227.

Mann:2014:Eb

[Man14b] Tony Mann. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):153, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.945138.

Mann:2015:BRR

[Man15a] Tony Mann. Book review:

Robert Recorde: the Life and Times of a Tudor Mathematician, by Gareth Roberts and Fenny Smith (editors). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):89, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.961703.

Mann:2015:END

[Man15b] Tony Mann. Editorial [Note on the deaths of Jackie Stedall and Ivor Grattan-Guinness]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):1, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.998955.

Mann:2015:ESJ

[Man15c] Tony Mann. Editorial: [Sixth Joint Meeting of the BSHM and the Canadian Society for the History and Philosophy of Mathematics]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):177, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1089041.

Held in Washington, DC, August 5–8, 2015.

Mann:2015:EPI

[Man15d] Tony Mann. Editorial [The passing of Ivor Grattan-Guinness and Jackie Stedall]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):93, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1036340.

Mann:2016:E

[Man16a] Tony Mann. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):89, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1180919.

Mann:2016:EJE

[Man16b] Tony Mann. Editorial [John Earle]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):159, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1237148.

Mann:2017:EME

[Man17] Tony Mann. Editorial [mathematics emerging: a tribute to Jakie Stedall and her influence on the history of mathematics]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):1, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1262312.

Meeting held at the Queen’s College, Oxford, April 9–10, 2016.

Mansfield:2023:MSR

[Man23] Daniel F. Mansfield. Mesopotamian square root approximation by a sequence of rectangles. British Journal for the History of Mathematics, 38(3):175–188, 2023. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2215652.

Marr:2006:MMT

[Mar06]  Alexander Marr.  Mak-

ing a mathematical textbook: Mutio Oddi’s Dello squadro. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(1):34–39, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600578019.

Martin:2014:NN

[Mar14] David Martin. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(2):146–152, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.894741.

Mason:2007:BPA

[Mas07] John Mason. Bartering problems in arithmetic books 1450– 1890. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):160–181, 2007. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430701581037.

Mann:2009:R

[MCC^{+}09] Tony Mann, Tony Crilly, Shelley Costa, Charles Care, Robin Wilson, and Jackie Stedall. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):191–199, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903088575.

McCartney:2011:PLJ

[McC11] Mark McCartney. The poetic life of James Clerk Maxwell. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):29–43, 2011.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903516138.

McCartney:2016:RPQ

[McC16] Mark McCartney. Research in progress The Queen’s College, Oxford, 27 February 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):249, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1215867.

McCallum:2018:MMR

[McC18]  Katie McCallum.  Mathe-

matics, manifest: a review of Mathematics: the Winton Gallery at the Science Museum. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):50–53, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1384201.

McDonald:2018:WBN

[McD18] Liam McDonald. Book review: Wonders beyond numbers: A brief history of all things mathematical, by Johnny Ball. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):201–203, 2018.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1504404.

McQuoid:2020:BRL

[McQ20] Annie McQuoid. Book review: Leonhard Euler’s Letters to a German Princess: A milestone in the history of physics textbooks and more, Ronald S. Calinger, Ekaterina (Katya) Denisova, and Elena N. Polyakhova, Morgan & Claypool, 2019, xvii + 214 pp, pb £78.95, hb £98.95, ISBN 978-1-64327-189-7. British Journal for the History of Mathematics, 35(3):251–253, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1797342.

Massa-Esteve:2017:MMI

[ME17] M. Rosa Massa-Esteve. Mengoli’s mathematical ideas in Leibniz’s excerpts. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1): 40–60, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/17498430.2016.
                                        1239807.

Morrison:2022:WPM

[MF22] Amie Morrison and Isobel Falconer. Women’s participation in mathematics in Scotland, 1730–1850. British Journal for the History of Mathematics, 37(1):2–23, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2053370.

Morris:2011:RFC

[MG11] Stephen Morris and Orlena Gotel. The role of flow charts in the early automation of applied mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(1):44–52, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903449207.

Marr:2010:R

[MGW^{+}10] Alexander Marr, Jeremy Gray, Benjamin Wardhaugh, Birgit Bergmann, and Jackie Stedall. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):181–189, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.497297.

Mason:2011:R

[MH11] John Mason and Keith Hannabuss.

Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):130–134, 2011.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.559896.

Mintz:2009:HLC

[Min09] Daniel Mintz. The hunt for the lost cities of Ptolemy. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):1–11, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802551913.

Maidment:2019:MWI

[MM19] Alison Maidment and Mark McCartney. ‘A man who has infinite capacity for making things go’: Sir Edmund Taylor Whittaker (1873– 1956). British Journal for the History of Mathematics, 34(3):179–193, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1619410.

Mazzotti:2009:R

[MMP^{+}09] Massimo Mazzotti, Mark Mc-Cartney, Anthony V. Piccolino, Snezana Lawrence, and Amirouche Moktefi. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):121–130, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430902848789.

Moktefi:2018:DLC

[Mok18] Amirouche Moktefi. Did Lewis Carroll own a copy of George Boole’s Laws of thought? An argument from the sale catalogues. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):21–28, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1384200.

Monteferrante:2011:MM

[Mon11] Sandra Monteferrante. Maya mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):71–79, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.504962.

Martins:2022:FIE

[MS22] Ana Patrícia Martins and Teresa Sousa. Formulations of the inclusion–exclusion principle from Legendre to Poincaré, with emphasis on Daniel Augusto da Silva. British Journal for the History of Mathematics, 37(3):212–229, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2082158.

Muntersbjorn:2012:IHH

[Mun12] Madeline Muntersbjorn. On the intellectual heritage of Henri Poincaré. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27 (2):107–118, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.658538.

Murphy:2016:GBW

[Mur16]  Donal Murphy.  George

Boole and Walsh’s delusions. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31 (2):123–127, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1113374.

Musto:2010:MT

[Mus10] Garrod Musto. Mathematical timelines. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):162–167, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498431003751393.

Nakane:2014:DWD

[Nak14] Michiyo Nakane. Did Weierstrass’s differential calculus have a limit-avoiding character? His definition of a limit in ϵ − δ style. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):51–59, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.831241.

Nasifoglu:2016:BRA

[Nas16] Yelda Nasifoglu. Book review: Architecture and mathematics from antiquity to the future, Volume I: Antiquity to the 1500s, edited by Kim Williams and Michael J. Ostwald/Architecture and mathematics from antiquity to the future, Volume II: The 1500s to the future, edited by Kim Williams and Michael J. Ostwald. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):261–263, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1208392.

Nyberg-Brodda:2021:BBN

[NB21] Carl-Fredrik Nyberg-Brodda.

The B B Newman spelling theorem. British Journal for the History of Mathematics, 36(2):117–131, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1904735.

Neumann:2008:HSA

[Neu08]  Peter M. Neumann.  The

history of symmetry and the asymmetry of history. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23 (3):169–177, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802304024.

NunoSilva:2023:LLF

[NF23] Jorge Nuno Silva and Pedro Jorge Freitas. The Loterias Lisbonenses of Francisco Giraldes Barba. British Journal for the History of Mathematics, 38(3):194–207, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2023.2258331.

Nossum:2015:SPS

[NK15] Rolf Nossum and Jan Kotůlek. The Society for the Protection of Science and Learning as a patron of refugee mathematicians. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):153–167, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1010778.

Nothaft:2020:MES

[Not20] C. Philipp E. Nothaft. Medieval Europe’s satanic ciphers: on the genesis of a modern myth. British Journal for the History of Mathematics, 35(2):107–136, 2020. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1726050.

Negrepontis:2016:TPI

[NT16] S. Negrepontis and G. Tassopoulos. Theodorus’ proofs of incommensurabilities with gnomons. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):15–30, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1055156.

Oaks:2022:ZNM

[Oak22] Jeffrey Oaks. Zero and nothing in medieval Arabic arithmetic. British Journal for the History of Mathematics, 37(3):179–211, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2115745.

Opitz:2012:R

[OHL12] Donald L. Opitz, Luke Hodgkin, and Snezana Lawrence. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(1):62–69, 2012. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.625522.

Opitz:2016:BRS

[Opi16]  Donald L. Opitz.  Book

review: Seduced by logic: Émilie Du Châtelet, Mary Somerville and the Newtonian revolution, by Robyn Arianrhod. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):156–158, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1170358.

Otero:2021:CG

[Ote21] Daniel E. Otero. Calculus gems. British Journal for the History of Mathematics, 36(3):219–221, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2021.1997199.

Padua:2017:PLB

[Pad17]  Sydney Padua.  Picturing

Lovelace, Babbage, and the Analytical Engine: a cartoonist in mathematical biography. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):214–220, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1315520.

Parshall:2006:BDT

[Par06] Karen Hunger Parshall. The British development of the theory of invariants (1841– 1895). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):186–199, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600964417.

Parshall:2017:PAA

[Par17] Karen Hunger Parshall. A plurality of algebras, 1200– 1600: algebraic Europe from Fibonacci to Clavius. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1): 2–16, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430.2016.
1225340.

Pisano:2016:NTD

[PB16] Raffaele Pisano and Paolo Bussotti. A Newtonian tale details on notes and proofs in Geneva edition of Newton’s Principia. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):160–178, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1183182.

Ponce-Campuzano:2015:VVC

[PCMA15] Juan Carlos Ponce-Campuzano and Miguel Ángel Maldonado-Aguilar. Vito Volterra’s construction of a nonconstant function with a bounded, non-Riemann integrable derivative. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):143–152, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1010771.

Peckhaus:2020:BRM

[Pec20] Volker Peckhaus. Book review: The mathematical world of Charles L. Dodgson (Lewis Carroll). British Journal for the History of Mathematics, 35(2):175–177, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1741781.

Periton:2015:MCT

[Per15] Cheryl Periton. The medieval counting table revisited: a brief introduction and description of its use during the early modern period. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):35–49, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.917392.

Petrocchi:2015:NTA

[Pet15] Alessandra Petrocchi. A new theoretical approach to sample problems and deductive reasoning in Sanskrit mathematical texts. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):2–19, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.976805.

Parker:2009:R

[PHN^{+}09] Rhian Parker, Ken Houston, Peter M. Neumann, Janet Beery, and David Klein. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1):50–61, 2009. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430802551962.

Pisano:2016:DMI

[Pis16] Raffaele Pisano. Details on the mathematical interplay between Leonardo da Vinci and Luca Pacioli. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):104–111, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1091969.

Peterson:2013:BRG

[PO13] Mark A. Peterson and Richard J. Oosterhoff. Between Raphael and Galileo: Mutio Oddi and the mathematical culture of late Renaissance Italy/ Galileo’s muse: Renaissance mathematics and the arts. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28 (2):107–111, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.761020.

Poole:2015:BRC

[Poo15] William Poole. Book review: The Correspondence of John Wallis (1616–1703) Volume IV (1672–April 1675), edited by Philip Beeley and Christoph J Scriba. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):174–175, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1037626.

Pope:2018:ADC

[PR18] Sue Pope and Leo Rogers.

Anglo–Danish conference on History of Mathematics in Education: Bath Spa, August 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):62–67, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1394713.

Pritchard:2006:BIG

[Pri06] Chris Pritchard. Bagatelle as the inspiration for Galton’s Quincunx. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):102–110, 2006. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600804118.

See comments [Ano07d, Pri08].

Pritchard:2008:MCC

[Pri08] Chris Pritchard. Mistakes concerning a chance encounter between Francis Galton and John Venn. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):103–108, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430801968068.

See [Pri06].

Przenioslo:2009:IMJ

[Prz09] Małgorzata Przenioslo. International mathematical journals published in Poland between the wars. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(1): 20–26, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

                                        /www.tandfonline.com/doi/
                                        full/10.1080/17498430802554487.

Rampling:2011:EME

[Ram11a] Jennifer M. Rampling. The Elizabethan mathematics of everything: John Dee’s ‘Mathematicall praeface’ to Euclid’s Elements. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):135–146, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.580136.

Ramsden:2011:BM

[Ram11b] Jenny Ramsden. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):128, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2011.560093.

Ranford:2021:BRH

[Ran21]  Paul Ranford.  Book re-

view: Hot molecules, cold electrons, Paul Nahin, Princeton University Press, 2020, 212 pp, $24.95, ISBN 978-0-6911-9172-0. British Journal for the History of Mathematics, 36(1):53–55, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1841477.

Rauff:2013:CWB

[Rau13] James Rauff. The chicken went into the bush and never came back: a note on infinity. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):97–100, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.727147.

Rauff:2016:AME

[Rau16] James V. Rauff. The algebra of marriage: an episode in applied group theory. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31 (3):230–244, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1222478.

Rice:2016:CCHa

[RB16a] Adrian Rice and Ezra Brown. Commutativity and collinearity: a historical case study of the interconnection of mathematical ideas. Part I. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):1–14, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1046037.

Rice:2016:CCHb

[RB16b] Adrian Rice and Ezra Brown. Commutativity and collinearity: a historical case study of the interconnection of mathematical ideas. Part II. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):90–103, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2015.1046038.

Recio:2021:KDB

[Rec21]  Gonzalo Luis Recio.  Ke-

pler’s derivation of the bisection of the earth’s orbit in Astronomia Nova. British Journal for the History of Mathematics, 36(3):151–169, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2021.1984061.

Reichenberger:2023:EHH

[Rei23] Andrea Reichenberger. Elli Heesch, Heinrich Heesch and Hilbert’s eighteenth problem: collaborative research between philosophy, mathematics and application. British Journal for the History of Mathematics, 38(3):208–228, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2023.2297522.

Rekvenyi:2019:PES

[Rek19]  Kamilla Rekvenyi.  Paul

Erdős’s mathematics as a social activity. British Journal for the History of Mathematics, 34(2):134–142, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1593036.

Rice:2006:BM

[Ric06] Adrian Rice. British mathematics 1837–1901. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21 (3):164–177, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430600964243.

Rice:2014:BRJ

[Ric14] Adrian Rice. Book review: James Clerk Maxwell. Perspectives on his life and work, by Raymond Flood, Mark Mc-Cartney and Andrew Whitaker (eds). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):219–221, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.919433.

Rice:2015:BGC

[Ric15a] Adrian Rice. Bshm/gresham college joint meeting on the history of statistics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):79–80, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.995344.

Rice:2015:OIG

[Ric15b]  Adrian Rice.  Obituary:

Ivor Grattan-Guinness (23 June 1941–12 December 2014). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):94–101, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1036338.

Rice:2016:BRT

[Ric16] Adrian Rice. Book review:

Taming the unknown: a history of algebra from antiquity to the early twentieth century, by Victor J Katz and Karen Hunger Parshall. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):81–83, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1080441.

Richardson:2019:WIW

[Ric19] John T. E. Richardson. Who introduced Western mathematicians to Latin squares? British Journal for the History of Mathematics, 34 (2):95–103, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1587998.

Ross:2019:DCT

[RK19] Helen Elizabeth Ross and Betty Irene Knott. Dicuil (9th century) on triangular and square numbers. British Journal for the History of Mathematics, 34(2):79–94, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1598687.

Robson:2015:SEM

[Rob15]  Eleanor Robson.  Subvert-

ing expectations: memories of editing with Jackie. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30 (3):178–182, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1055902.

Robertson:2018:BRG

[Rob18] Edmund Robertson. Book review: Giovanni Domenico Cassini: A Modern Astronomer in the 17th Century, by Gabriella Bernardi. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33 (3):191–192, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1450034.

Rodhe:2017:FBG

[Rod17a] Staffan Rodhe. A forgotten booklet by Goldbach now rediscovered and three versions of its contents. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):75–90, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1225628.

Rodriguez:2017:FRB

[Rod17b] L. Rodríguez. Frigyes Riesz between the two World Wars. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (3):235–245, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1326216.

Rodal:2018:LAE

[Rod18]  Jocelyn Rodal.  Book re-

view: Literature after Euclid: the geometric imagination in the long Scottish Enlightenment, by Matthew Wickman. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):55–57, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1308122.

Rogers:2020:BRH

[Rog20] Leo Rogers. Book review:

The history of mathematics: a source-based approach volume 1. British Journal for the History of Mathematics, 35(2):171–173, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1741783.

Rowlett:2014:BBL

[Row14]  Peter Rowlett.  Brass for

brain: Lord Kelvin and tide prediction. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):211–213, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.916114.

Rowlett:2015:UIM

[Row15] Peter Rowlett. ‘The unplanned impact of mathematics’ and its implications for research funding: a discussion-led educational activity. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):67–74, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.945136.

Rowlett:2024:GAA

[Row24] Peter Rowlett. Generative AI and accuracy in the history of mathematics. British Journal for the History of Mathematics, 39(1):64–69, 2024. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2024.2312789.

Romero-Vallhonesta:2018:MSA

[RVME18] Fàtima Romero-Vallhonesta and M. Rosa Massa-Esteve.

The main sources for the Arte Mayor in sixteenth century Spain. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):73–95, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1419704.

Satterley:2021:MBF

[Sat21] Renae Satterley. Mathematical books and Frankfurt book fair catalogues: the acquisition of mathematical works by Robert Ashley in early modern London. British Journal for the History of Mathematics, 36(2):95–116, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2021.1891804.

Schwermer:2015:BRE

[Sch15] Joachim Schwermer. Book review: Emil Artin’s Iceland Journal 1925: “A World of Good”, edited by Tom Artin (ed) and Karin Tate (trans). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):172–173, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1036337.

Schubring:2021:BRD

[Sch21] Gert Schubring. Book review: Dirichlet. A mathematical biography, Uta C. Merzbach, Birkhäuser, 2018, xix + 311 pp., hardback £89.99, ebook £71.50, ISBN 978-3-03001-071-3. British Journal for the History of Mathematics, 36(1):55–58, 2021. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2020.1823606.

Schneider:2022:WSJ

[Sch22] Martina R. Schneider. ‘What to solve?’ — on Judita Cofman’s research on mathematics and its teaching. British Journal for the History of Mathematics, 37(1):96–98, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2047575.

Segre:2020:TSW

[Seg20] Michael Segre. Tartaglia’s science of weights and mechanics in the sixteenth century. Selections from Quesiti et inventioni diverse: Books VII–VIII. British Journal for the History of Mathematics, 35(1):84–85, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1678829.

Seltman:2008:AAP

[Sel08] Muriel Seltman. The Artis analyticae praxis of Harriot and Warner in focus. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):73–80, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430801968043.

Senn:2016:BRC

[Sen16] Stephen Senn. Book review: A Chronicle of Permutation Statistical Methods: 1920–2000, and Beyond, by K. J. Berry, J. E. Johnston, and P. J. W. Mielke. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(2):155–156, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1129087.

Salas-Garcia:2022:IEH

[SGPBGL22] Irene Salas-García, Irene Polo-Blanco, and María José González-López. Instrument for evaluating historical resources in mathematics textbooks. British Journal for the History of Mathematics, 37(3):238–257, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2101799.

Sharp:2016:FRP

[Sha16] John Sharp. Folding the regular pentagon. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):179–188, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1162402.

Sialaros:2015:BRK

[Sia15] Michalis Sialaros. Book review: The King of Infinite Space: Euclid and his Elements, by David Berlinski. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):83–85, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.928840.

Silva:2011:MG

[Sil11] Jorge Nuno Silva. On mathematical games. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26 (2):80–104, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.560511.

Simmons:2008:WRH

[Sim08] Charlotte Simmons. William Rowan Hamilton and George Boole. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):96–102, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430801968050.

Simpson:2014:JMJ

[Sim14] Richard Simpson. A joint meeting of the James Clerk Maxwell Foundation and the British Society for the History of Mathematics in celebration of the 400th anniversary of the publication of John Napier’s Mirifici Logarithmorum Canonis Descriptio. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):213–214, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.948348.

Simpson:2018:BRF

[Sim18] Richard Simpson. Book review: Finding Fibonacci: the quest to rediscover the forgotten mathematical genius who changed the world, by Keith Devlin. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):136–137, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1396782.

Stalpers:2018:EKK

[SK18] Lukas J. A. Stalpers and Edward L. Kaplan. Edward L. Kaplan and the Kaplan–Meier Survival Curve. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33 (2):109–135, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1450055.

Smith:2008:IAM

[Smi08]  Fenny Smith.  The influ-

ence of Amatino Manucci and Luca Pacioli. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(3):143–156, 2008. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802304016.

Smith:2015:RP

[Smi15] Fenny Smith. Research in progress. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):169–171, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2015.1035572.

Sonar:2014:BRH

[Son14] Thomas Sonar. Book review: The history of mathematics: a very short introduction, by Jacqueline Stedall. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29 (3):217–219, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.920217.

Sonar:2017:LDN

[Son17] Thomas Sonar. … in the darkest night that is … Briggs, Blundeville, Wright, and the misconception of finding latitude. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):20–29, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1225337.

Sorensen:2013:WAA

[Sør13] Henrik Kragh Sørensen. What’s Abelian about abelian groups? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1):37–51, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2013.729919.

Sorensen:2017:SAP

[Sør17] Henrik Kragh Sørensen. Studying appropriations of past lives: using metabiographical approaches in the history of mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):186–197, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1355641.

Spencer:2019:BRP

[Spe19] Fiona Spencer. Book review: The path to post-Galilean epistemology: reinterpreting the birth of modern science, by Danilo Capecchi. British Journal for the History of Mathematics, 34(1):66–68, 2019. CODEN ???? ISSN 2637-

                                        5494.  URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1543994.

Sharp:2009:BM

[SPR09] John Sharp, Martin Perkins, and Jenny Ramsden. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):114–116, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430902848649.

Steiner:2013:R

[SR13] Katherine Steiner and Leo Rogers. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(1): 54–58, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/
full/10.1080/17498430.2013.
736276.

Siegmund-Schultze:2022:ALT

[SS22] Reinhard Siegmund-Schultze.

Another look at the two Egyptian pyramid volume ‘formulas’ of 1850 BCE. British Journal for the History of Mathematics, 37(3):171–178, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2106061.

Stamhuis:2015:BRS

[Sta15] Ida Stamhuis. Book review:

Statistics, Public Debate and the State, 1800–1945: A Social, Political and Intellectual History of Numbers, by Jean-Guy Prévost and Jean-Pierre Beaud. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):87–88, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.956275.

Stedall:2006:ECV

[Ste06] Jackie Stedall. Editorial [Conference on Victorian Mathematics]. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(3):163, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600963633.

Held in Oxford, July 2006.

Stedall:2008:E

[Ste08]  Jackie Stedall.  Editorial.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(2):57, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430802062341.

Stephenson:2009:GDL

[Ste09] Craig Stephenson. George Darwin’s lectures on Hill’s lunar theory. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(3):159–171, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903008409.

Stedall:2010:Ea

[Ste10a]  Jackie Stedall.  Editorial.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):1, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903423459.

Stedall:2010:Eb

[Ste10b]  Jackie Stedall.  Editorial.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):57, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498431003696259.

Stedall:2010:Ec

[Ste10c]  Jackie Stedall.  Editorial.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(3):131, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2010.497299.

Stedall:2013:E

[Ste13]  Jackie Stedall.  Editorial.

BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):65, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.755074.

Stedall:2014:NN

[Ste14] Jackie Stedall. News and notices. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(1):80–81, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.855973.

Stenhouse:2019:NHM

[Ste19]  Brigitte Stenhouse.  29th

Novembertagung on the history of mathematics. British Journal for the History of Mathematics, 34(2):133, 2019. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2019.1589782.

Stenhouse:2023:BMN

[Ste23] Brigitte Stenhouse. BSHM meeting news. British Journal for the History of Mathematics, 38(3):231–232, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2023.2271815.

Stedall:2012:BM

[SW12] Jackie Stedall and Robin Wilson. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(2):126, 2012. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.676964.

Skiena:2017:WBW

[SW17] Steven S. Skiena and Charles B.

Ward. Who’s bigger? Where computer scientists really rank. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32 (3):257–264, 2017. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1317979.

Takei:2017:RPS

[Tak17] Masato Takei. A remark on the paper ‘Sieve of war: the legacy of Jitsuro Nagura’ by Daniel Tisdale. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):170, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1281030.

See [Tis13].

Tanswell:2020:BRV

[Tan20] Fenner Stanley Tanswell. Book review: 99 Variations on a proof. British Journal for the History of Mathematics, 35(2):173–175, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2020.1735618.

Taylor:2011:VGB

[Tay11] Katie Taylor. Vernacular geometry: between the senses and reason. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(3):147–159, 2011. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.580137.

Teather:2019:CDF

[TCP19] Anne Teather, Andrew Chamberlain, and Mike Parker Pearson. The chalk drums from Folkton and Lavant: Measuring devices from the time of Stonehenge. British Journal for the History of Mathematics, 34(1):1–11, 2019. CO-DEN ???? ISSN 2637-

                                        5494.  URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2018.1555927.

Thanailaki:2010:BSB

[Tha10]  Polly Thanailaki.  Break-

ing social barriers: Florentia Fountoukli (1869–1915). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):32–38, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903321166.

Tisdale:2013:SWL

[Tis13]  Daniel Tisdale.  Sieve of

war: the legacy of Jitsuro Nagura. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(3):143–155, 2013. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.793949.

See remark [Tak17].

Taub:2010:R

[TMIvM10] Liba Taub, Daniel V. Mintz, Annette Imhausen, and Jan van Maanen. Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):44–51, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903407924.

Tou:2019:BIL

[Tou19]  Erik R. Tou.  Bernoul-

lian influences on Leonhard Euler’s early fluid mechanics. British Journal for the History of Mathematics, 34(2):104–117, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1573557.

Turner:2015:PMB

[Tur15]  Colin Turner.  Physicists

& mathematicians of Belfast

— joint BSHM/Institute of Physics event. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):77, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.950481.

VanBrummelen:2019:BRJ

[Van19] Glen Van Brummelen. Book review: Jacqueline Feke: Ptolemy’s philosophy: mathematics as a way of life. British Journal for the History of Mathematics, 34 (2):118–119, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1582259.

vanBrummelen:2010:FSB

[vB10] Glen van Brummelen. Filling in the short blanks: musings on bringing the historiography of mathematics to the classroom. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):2–9, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430903321125.

Verburgt:2014:RIE

[Ver14] Lukas M. Verburgt. Remarks on the idealist and empiricist interpretation of frequentism: Robert Leslie Ellis versus John Venn. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):184–195, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.889269.

Verburgt:2016:RLE

[Ver16] Lukas M. Verburgt. Robert Leslie Ellis, William Whewell and Kant: the role of Rev H F C Logan. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(1):47–51, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1035582.

Verburgt:2018:LRL

[Ver18] Lukas M. Verburgt. A letter of Robert Leslie Ellis to William Walton on probability. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):96–108, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1437244.

vanMaanen:2006:DMR

[vM06] Jan van Maanen. Diagrams and mathematical reasoning: some points, lines, and figures. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 21(2):97–101, 2006. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430600803532.

vanWeerden:2022:AQA

[vW22]  Anne van Weerden.  Al-

ice without quaternions: another look at the mad teaparty. British Journal for the History of Mathematics, 37(3):230–237, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2085446.

vanWeerden:2018:MGA

[vWW18] Anne van Weerden and Steven Wepster. A most gossiped about genius: Sir William Rowan Hamilton. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):2–20, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1400821.

Wagner:2018:BRM

[Wag18] Roy Wagner. Book review: Mathematical knowledge and the interplay of practices, by José Ferreirós. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):141–143, 2018.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1388101.

Walker:2014:BRP

[Wal14] Mairi Walker. BSHM research in progress. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):210–211, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.917390.

Wardhaugh:2007:PRM

[War07] Benjamin Wardhaugh. Poor Robin and Merry Andrew: mathematical humour in Restoration England. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 22(3):151–159, 2007. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701581011.

Wardhaugh:2009:OM

[War09] Benjamin Wardhaugh. Other meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 24(2):117–118, 2009. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430902848656.

Wardhaugh:2010:LUP

[War10] Benjamin Wardhaugh. ‘Let us put on the shade of Newton’: Isaac Newton on stage, 1829–2006. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):67–80, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498431003659844.

Wardhaugh:2011:OM

[War11] Benjamin Wardhaugh. Other meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 26(2):129, 2011. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2011.561090.

Wardhaugh:2014:PME

[War14] Benjamin Wardhaugh. Printing mathematics in the early modern world research symposium. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):215–216, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.916112.

Wardhaugh:2017:CHO

[War17] Benjamin Wardhaugh. Charles Hutton: ‘one of the greatest mathematicians in Europe’? BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1):91–99, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1236319.

Wardhaugh:2018:Ea

[War18a] Benjamin Wardhaugh. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):1, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1400804.

Wardhaugh:2018:Eb

[War18b] Benjamin Wardhaugh. Editorial. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):147, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2018.1521595.

Wardhaugh:2022:E

[War22] Benjamin Wardhaugh. Editorial. British Journal for the History of Mathematics, 37(1):1, 2022. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.

                                        tandfonline.com/doi/full/
                                        10.1080/26375451.2022.2039512.

Weeks:2010:BM

[Wee10] Chris Weeks. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(1):39–40, 2010. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430903384065.

Wess:2012:AAM

[Wes12] Jane Wess. Avoiding arithmetic, or the material culture of not learning mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(2):82–106, 2012. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2012.622639.

Wess:2016:JWM

[Wes16]  Jane Wess.  John Wal-

lis (1616–1703). Mathematics, music theory, and cryptography in 17th century Oxford. Oxford University Mathematical Institute, 9 June 2016. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):252–253, 2016. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2016.1215868.

Wess:2018:BM

[Wes18] Jane Wess. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):70–72, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1394715.

Wess:2020:CMH

[Wes20] Jane Wess. Colin Maclaurin (1698–1746) and his contemporaries on wind and water: the local and the universal. British Journal for the History of Mathematics, 35(1):63–83, 2020. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1701861.

Wess:2022:STM

[Wes22] Jane Wess. Symbols and things: mathematics in the age of steam. British Journal for the History of Mathematics, 37(1):82–85, 2022. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2022.2036410.

Wess:2024:BRE

[Wes24] Jane Wess. Benjamin Robins:

elegant mathematics versus experimental inconvenience? British Journal for the History of Mathematics, 39(1):3–25, 2024. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2024.2317696.

Wilson:2021:BFF

[WF21] Robin Wilson and Raymond Flood. The BSHM: the first fifty years. British Journal for the History of Mathematics, 36(2):76–94, 2021. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2021.1919462.

Whiteside:2014:JNC

[Whi14] D. T. Whiteside. ‘And John Napier created logarithms… ’. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):154–166, 2014. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.874855.

Whitaker:2015:TIH

[Whi15] Andrew Whitaker. The Third Irish History of Mathematics Conference. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(3):260, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1057008.

Wilson:2008:ECM

[Wil08]  Robin Wilson.  Euler’s

combinatorial mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 23(1):13–23, 2008. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430701799175.

Wilson:2013:BM

[Wil13] Robin Wilson. BSHM meetings. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 28(2):101–106, 2013. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2013.760060.

Wilson:2014:FEM

[Wil14] Robin Wilson. In the footsteps of Euler and MacMahon: combinatorics, the mathematics that counts. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 29(3):196–209, 2014. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.916115.

Wilson:2015:BSF

[Wil15a] Robin Wilson. British Science Festival. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):78–79, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.966436.

Wilson:2015:BCM

[Wil15b] Robin Wilson. BSHM Christmas meeting. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(2):168–169, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.1001619.

Wilson:2015:HMF

[Wil15c] Robin Wilson. History of mathematics film day. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):79, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2014.993114.

Wilson:2015:HMS

[Wil15d]  Robin Wilson.  The his-

tory of mathematics in 300 stamps. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):80–81, 2015.

CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.993113.

Wilson:2015:JMB

[Wil15e] Robin Wilson. Joint meeting of the BSHM and the IMA North-West Branch. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):75, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.942821.

Wilson:2015:NWM

[Wil15f] Robin Wilson. Non-Western mathematics. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30(1):77–78, 2015. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2014.942822.

Wilson:2015:SBC

[Wil15g] Robin Wilson. Sixth BSHM– CSHPM Joint Meeting. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 30 (3):255–259, 2015. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2015.1086071.

Wilson:2017:GPGa

[Wil17a] Robin J. Wilson. The Gresham Professors of Geometry. Part 1: the first one hundred years. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):125–135, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1236317.

Wilson:2017:GPGb

[Wil17b] Robin J. Wilson. The Gresham Professors of Geometry. Part 2: the next three hundred years. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):136–148, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2016.1236318.

Wilson:2018:LEM

[Wil18a]  Robin Wilson.  Book re-

view: Leonhard Euler: mathematical genius in the Enlightenment, by Ronald S. Calinger. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):54–55, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1308123.

Wilson:2018:BLC

[Wil18b]  Robin Wilson.  BSHM/

Lewis Carroll Society: The mathematical world of C. L. Dodgson (Lewis Carroll). BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(1):69–70, 2018. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2017.1384202.

Weeks:2010:R

[WM10] Chris Weeks and Tony Mann.

Reviews. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 25(2):127–128, 2010. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498431003664471.

Wood:2019:BRR

[Woo19] Miranda Wood. Book review: Robin Wilson, Euler’s pioneering equation: the most beautiful theorem in mathematics. British Journal for the History of Mathematics, 34(2):120–123, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1582261.

Wong:2012:MEH

[WT12] Ngai Ying Wong and Kwok Chun Tang. Mathematics education in Hong Kong under colonial rule. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 27(2):119–125, 2012. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2012.628480.

Yilmaz:2017:ARW

[YO17] Zuhal Yilmaz and Suat Eren Ozyigit. Analysis of real world problems in mathematics textbooks of early twentieth and twenty-first century Turkish education: political and social reflections. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(2):171–182, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/17498430.2016.1247323.

Zelbo:2019:RMA

[Zel19]  Sian Zelbo.  The recre-

ational mathematics activities of ordinary nineteenth century Americans: a case study of two mathematics puzzle columns and their contributors. British Journal for the History of Mathematics, 34(3):155–178, 2019. CO-DEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

tronic). URL http://www.
tandfonline.com/doi/full/
10.1080/26375451.2019.1646522.